### v2

# What Matters to Enhance Traffic Rule Compliance of Imitation Learning for Automated Driving

Abstract
More research attention has recently been given to end-to-end autonomous driving technologies where the entire driving pipeline is replaced with a single neural network because of its simpler structure and faster inference time. Despite this appealing approach largely reducing the components in the driving pipeline, its simplicity also leads to interpretability problems and safety issues [1]. The trained policy is not always compliant with the traffic rules and it is also hard to discover the reason for the misbehavior because of the lack of intermediate outputs. Meanwhile, sensors are also critical to autonomous driving’s security and feasibility to perceive the surrounding environment under complex driving scenarios. In this paper, we proposed P-CSG, a penalty-based imitation learning approach with cross semantics generation sensor fusion technologies to increase the overall performance of end-to-end autonomous driving. In this method, we introduce three penalties - red light, stop sign, and curvature speed penalty to make the agent more sensitive to traffic rules. The proposed cross semantics generation helps to align the shared information from different input modalities. We assessed our model’s performance using the CARLA leaderboard - Town 05 Long benchmark and Longest6 Benchmark, achieving an impressive driving score improvement. Furthermore, we conducted robustness evaluations against adversarial attacks like FGSM and Dot attacks, revealing a substantial increase in robustness compared to baseline models. More detailed information, such as code base resources, and videos can be found at https://hk-zh.github.io/p-csg-plus.
Index Terms:
Imitation Learning, Multi-modality Sensor Fusion, End-to-end Autonomous Driving
I Introduction
End-to-end autonomous driving integrates the perception and decision layers into one deep neural network. Enhancing the policy learning algorithm and improving perception abilities are two crucial aspects that can significantly enhance the overall performance of autonomous driving.
Imitation learning [2] and reinforcement learning [3] are two main learning paradigms unutilized end-to-end autonomous driving systems for the decision-making part. Although both approaches have been successful in the field of autonomous driving, they each have their own drawbacks. Imitation learning only learns the behaviors limited to the training data. This means that the model may struggle to generalize to new and unfamiliar situations. The quality of the expert policy also has a significant impact on the performance of imitation learning. The design of the reward function plays a critical role in the effectiveness of reinforcement learning for autonomous driving systems, and creating an appropriate reward function can be a complex task. Additionally, there is a risk that reinforcement learning models may learn unexpected and potentially dangerous behaviors in such safety-critical systems. In this article, we propose a penalty-based imitation learning paradigm that introduces the "penalty"/"reward" concept from reinforcement learing into imitation learning and aims to 1) achieve higher learning efficiency compared to reinforcement learning (learning from the expert demonstration) and 2) make the agent more sensitive to the traffic rules by involving some penalties designed in a simple manner.
The perception component in the neural network is used to extract essential information about the surrounding environment. Despite some Lidar-based approaches [4] [5], which leverages Lidar sensor input and HD maps, demonstrate impressive performance, relying on HD maps is not a feasible option as it requires substantial resources to create and maintain. They may not be universally available for all areas and regions. For this reason, an increasing number of research [6] [7] [8] [9] recently concentrate on multi-modality sensors (e.g Lidar and Camera) based approaches. Sensor fusion technologies are widely used in 2D object and 3D object detection. Classified by the types of point cloud representation of Lidar input, there exist three main approaches, namely point-based [10] [11], voxel-based [12] [13] [14] [15], and range-view-based [16] [17] [18] [19] approaches. More recently, with the popularity of attention mechanisms, many researchers are trying to use Transformer [20] to integrate multimodal information [7] [21] [8].In this paper, We shift our approach by extracting and aligning shared information from diverse modalities, enabling the model to extract the global context within multi-modalities
We propose Penalty-based Imitation Learning with Cross Semantics Generation (P-CSG). We evaluate our model on Carla Leaderboard - Town 05 Long Benchmark and Longest6 Benchmark. It achieves state-of-the-art performance, especially in collision and traffic rule violations. We also analyze the performance and robustness of our autonomous driving model under two types of adversarial attacks: FGSM and Dot attacks. The results show that our model outperforms the other baselines from the perspective of safety.
Our main contributions to this paper can be summarized as follows:
  • Our innovative multi-sensor fusion technique involves aligning shared information from various modalities, thus simplifying the process of extracting the global context across diverse modalities.
  • We proposed a penalty-based imitation learning approach that leverages constraint optimizations to make the imitation learning model more sensitive to traffic rule violations.
II Related Works
There are currently two primary methods used in autonomous driving technology: modular and end-to-end approaches. The modular approach involves a detailed series of software modules that work together to control the vehicle. By employing a modular approach, engineering teams can focus on specific and well-defined subtasks and make enhancements independently throughout the entire system. This allows the system to remain operational as long as the intermediate outputs remain functional. On the other hand, the end-to-end approach considers the entire driving process as one comprehensive learning task. It leverages the ability of the neural network to learn task-specific features and perform task-specific actions. Contrary to the modular approach, the agent learns without relying on explicit rules or hand-engineered features, namely the human-defined information bottlenecks. This means that the vehicle can adapt to different driving scenarios and environments and potentially achieve better performance than systems that rely on traditional rule-based approaches. End-to-end approaches have shown great success in computer vision, such as object detection, object tracking, and semantic segmentation. The success of these tasks builds a solid foundation for end-to-end autonomous driving. It is reasonable to assume that in the future, the end-to-end approach will also have the ability to operate a vehicle autonomously. Lately, there has been a growing trend in research to utilize behavior cloning for training agents to accomplish autonomous driving tasks. This involves the agent learning from demonstrations provided by experts in the field [7] [8] [6] [9]. They collect the demonstration in CARLA [22], a simulation urban environment in which the researchers and developers can test and evaluate their algorithms and models for autonomous driving. They employ strategies like enhancing multi-modality sensor fusion [7] [8], introducing safety module [8] to avoid dangerous behavior and improve the overall performance of autonomous driving. In our approach, penalties for traffic rules are introduced to ensure the safety of driving. We also introduce the cross semantics generation, which generates the semantic information of the other modality, to capture the global context across different modalities.
III Methodologies
In this section, we propose a multi-sensor fusion approach and a penalty-based imitation learning method for end-to-end autonomous driving.
Feature ExtractionCross Semantic GenerationAuxiliaryPenalty-based Imitation LearningRefer to captionRefer to captionRefer to captionptResnet50ptResnet18pt512pt512ptDecoderRefer to captionRefer to captionptMLPptMLPContractive Losspt128Refer to captionRefer to captionRefer to captionpt4Measurementspt1156Refer to captionRed Light PenaltyRefer to captionStop Sign PenaltyRefer to captionSpeed PenaltyptMLPpt64ptGRUptptGoalLocationptptptWaypointsptPIDRefer to captionRefer to captionPenalties
Figure 1: This is an overview of our proposed Penalty-based Imitation Learning with Cross Semantics Generation. Let’s start with the first feature extraction part. The top-down LiDAR pseudo image is fed into a ResNet18 architecture to extract the features from the LiDAR input. The front camera image is fed into a ResNet50 architecture to extract the features from the camera input. Then, the Cross Semantic Generation is conducted to cross-generate the front and top-down semantic segmentations in the frame of the front image and LiDAR pseudo image. We apply the contrastive loss the align the shared features of the camera and LiDAR inputs. The resampled shared embedding from latent space will be used to generate the waypoints. The features extracted from the front camera image are also leveraged to perform auxiliary tasks ’stop sign indication’ and ’red light indication’. In terms of penalty-based imitation learning, the LiDAR features, camera features, shared features, and measurements are concatenated and fed into an MLP network to reduce the dimension. A GRU structure is followed to recurrently generate the waypoints, also taking the goal location as input. The generated waypoints will be applied with the three penalties, namely, red light penalty, stop sign penalty, and speed penalty. The task of the PID controller is to map the waypoints into the action of steer, throttle, and brake.
III-A Problem Setting
The task we concentrate on is point-to-point navigation in an urban setting where the goal is to complete a route with safe reactions to dynamic agents such as moving vehicles and pedestrians. The traffic rules like red lights and stop signs should be followed. We consider the problem of learning a goal-conditioned policy $\pi(\bm{a}|\bm{s},\bm{g})$ that outputs action $\bm{a}\in\mathcal{A}$, conditioned on the current observed state $\bm{s}\in\mathcal{S}$ and a high-level goal location $\bm{g}$ provided by GPS. The environment we used to train and test the performance of the trained agent is CARLA. The specific environment setting we used can be characterized by the following statements:
  • A state space $\mathcal{S}\subset\{(\mathbb{R}^{H_{\textrm{c}}\times W_{\textrm{c}}\times 3},\mathbb{R}^{H_{\textrm{L}}\times W_{\textrm{L}}\times 2},\mathbb{R}^{4})\}$ which is the combination of the observation of camera with the shape of $H_{\textrm{c}}\times W_{\textrm{c}}\times 3$, bird’s eye view LiDAR input with the shape of $H_{\textrm{L}}\times W_{\textrm{L}}\times 2$, and measurements of current throttle, brake, steer, and velocity.
  • A multidimensional action space $\mathcal{A}\subset\mathbb{R}^{3}$. The action space contains the parameters of steer, throttle, and brake to drive the agent to finish tasks.
  • A goal space consists of high-level goal location provided as GPS coordinates.
III-A1 Observation Space
Similar to prior works [23], [24] [7] [21], the LiDAR point cloud data is transformed into a 2-bin histogram by mapping it onto a 2D BEV grid of fixed resolution. To encompass a 32m x 32m BEV grid, we limit our analysis to points located within 32m in front of the ego vehicle and 16m on each side. We partition the grid into blocks of 0.125m $\times$ 0.125m, which gives us a resolution of 256 $\times$ 256 pixels. The height dimension is discretized into two bins, distinguishing points that are either on/below or above the ground plane. This results in a two-channel pseudo image input $\mathbb{R}^{256\times 256\times 2}$.
In terms of RGB input, we leverage three cameras positioned to face forward, 60 degrees to the left, and 60 degrees to the right. These cameras have a horizontal field of view (FOV) of 120 degrees. The image from the cameras has a resolution of $400\times 300$ pixels. The forward image is cropped to $300\times 160$ pixels and the 60 degrees to the left and right images are cropped to $160\times 234$ pixels. we want to use cropping to eliminate radial distortion at the edges. We combine these three cropped images into a single image input with a resolution of $768\times 160$ pixels.
In addition, the current throttle, brake, steering, and velocity measurements are included in the observation space to provide a point of reference for the agent.
III-A2 Action Space
Instead of directly predicting the next action of throttle, steer, and brake, we estimate the future waypoints $\mathcal{W}$ of the ego-vehicle in bird’s eye view space which is centered at the ego vehicle’s current coordinate frame. The trajectory consists of a sequence of 2D waypoints $\mathcal{W}=\{\bm{w_{t}}=(x_{t},y_{t})\}^{T}_{t=1}$. In our settings, $T=4$ which means the agent needs to predict 4 waypoints in the future for each frame. Then we use two fine-tuned PID controllers for lateral and longitudinal control to obtain steer, throttle, and brake values from the predicted waypoints. These two PID controllers are not involved in the training process.
III-A3 Goal Location
According to CARLA [22] 0.9.10’s protocol, the high-level goal locations $\bm{g}$ are provided as GPS coordinates. This goal location $\bm{g}$ is sparse (hundreds of meters apart) which can only be used as guidance. We kindly remind the readers that the goal location is different from those predicted waypoints, which are dense, and only a few meters away from each other.
III-B Policy Learning
In this paper, we plan to use the imitation learning mechanism to learn the goal-conditioned policy $\pi(\bm{a}|\bm{s},\bm{g})$. As we discussed previously, we intend to predict the future waypoints instead of directly learning the actions. An expert policy is applied in the environment to collect a dataset $\mathcal{D}=\{(\bm{s},\mathcal{W})\}^{N}$ with the size of $N$, which contains multiple tuples of the current observation $\bm{s}$ and a set of waypoints $\mathcal{W}$ in the future timesteps. Similar to previous works [7], [25], we also use $l_{1}$-norm based loss function because of its demonstrated robustness to outliers. For each input, the policy learning loss can be formalized as:
$\mathcal{L}_{\textrm{pl}}=\sum_{t=1}^{T}||\bm{\hat{w}_{t}}-\bm{w_{t}}||_{1}$ (1)
where $\bm{\hat{w}_{t}}$ is the $t$-th predicted waypoint and $\bm{w_{t}}$ is the t-th ground truth waypoint produced by the expert policy. $\mathcal{L}_{\textrm{pl}}$ is the loss function of policy learning.
III-C Cross Semantics Generation
The motivation for our approach is based on the fact that multi-modality inputs have common information and also their own unique information. For instance, the vehicle’s and pedestrian’s shape and location are the shared information of LiDAR and Camera input. The term unique information refers to information that is not present in other sources of input. Extracting and aligning the common information from both modalities, helps the decision network to find the logical connection between modalities and capture their global context since the extracted features are more reasonable and easier to understand. There are two problems we want to solve in this section. They are 1) How to find the shared information of these two modalities? 2) How to align the extracted shared information in the same space?
To address the first problem, we propose the cross semantics generation approach. We plan to utilize the LiDAR input to produce semantic information from the camera and also use the camera input to generate semantic information from the LiDAR. As Figure 1 demonstrates, the information flows for the LiDAR and camera semantic information generation are crossed since we used the information from one modality to generate the semantic segmentation of the other modality. That is the reason we name our approach cross semantics generation. The RGB image and the bird’s eye view of the LiDAR pseudo image are fed into two residual networks [26] to extract the features of these two modalities. We utilize two linear layers to extract the shared information of the camera and LiDAR. Those from the camera input extracted features are used to generate the top-down semantic segmentation which is aligned with the LiDAR input and those from the LiDAR input extracted features are used to generate the semantic segmentation of the front image. The reason we used the information to generate the semantic segmentation is that 1) semantic segmentation has less noise compared to the original data and 2) with the semantic segmentation we can filter out the unique information such as traffic light color and traffic sign pattern. In our setup, the semantic segmentation contains 4 channels, namely the drivable area, the non-drivable area, the object in the drivable areas (vehicle and pedestrians), and others.
We define $y_{\textrm{front}}$ as the ground truth front segmenation tensor with the shape of $160\times 768\times 4$ and $\hat{y}_{\textrm{front}}$ as the output of the front view decoder with the same shape. We also define $y_{\textrm{td}}$ as the ground truth top-down segmentation tensor with the shape of $256\times 256\times 4$ and $\hat{y}_{\textrm{td}}$ as the output of the top-down view decoder with the same shape. The reconstruction losses can be written as $\mathcal{L}_{\textrm{front}}=\mathcal{L}_{\textrm{CE}}(\hat{y}_{\textrm{front}},y_{\textrm{front}})$ and $\mathcal{L}_{\textrm{td}}=\mathcal{L}_{\textrm{CE}}(\hat{y}_{\textrm{td}},y_{\textrm{td}})$, where $\mathcal{L}_{\textrm{CE}}(\cdot)$ indicates the cross entropy loss. $\mathcal{L}_{\textrm{front}}$ and $\mathcal{L}_{\textrm{td}}$ represent the loss function of front reconstruction loss and top-down reconstruction loss, respectively.
To align the extracted shared information into the same space, we leverage the contrastive loss. The primary goal of contrastive loss is to learn an embedding space where similar examples are placed closer together, while dissimilar examples are placed farther apart. This is achieved by defining a distance metric between the embeddings and then minimizing the distance between similar examples while maximizing the distance between dissimilar examples. Note that the embeddings extracted from LiDAR and camera inputs are mapped into two Gaussian distributions by generating mean and variance, using MLP networks. Suppose $P_{1}$ and $P_{2}$ represent two batches of Gaussian distributions, derived from RGB and LiDAR inputs. Each batch contains $N_{b}$ number of distributions. The contrastive loss can be formulated as follows:
$\mathcal{L}_{\textrm{align}}({P_{1}},{P_{2}})=\bm{E}\odot\bm{D}+(1-\bm{E})\odot\max\{0,\epsilon_{\textrm{a}}-\bm{D}\}$ (2)
where $\bm{D}$ is the matrix consisting of $d_{ij}=d(p_{1}^{i},p_{2}^{j})$. $p_{1}^{i}$ is the $i$-th distribution in ${P_{1}}$ and $p_{2}^{j}$ is the $j$-th distribution in ${P_{2}}$. $d(p_{1}^{i},p_{2}^{j})$ indicates the metric to measure the difference of two distributions. $\bm{E}$ is the matrix consisting of $e_{ij}$ which indicates if $p_{1}^{i}$ and $p_{2}^{j}$ belong to the same frame. $\epsilon_{\textrm{a}}$ is the threshold difference between similar samples and dissimilar samples. $\odot$ represents the Hadamard product of two matrices. In our setting, we use the symmetric version of Kullback–Leibler divergence
$d(p_{1}^{i},p_{2}^{j})=\frac{1}{2}\textrm{KL}(p_{1}^{i}||p_{2}^{j})+\frac{1}{2}\textrm{KL}(p_{2}^{j}||p_{1}^{i})$ (3)
where $\textrm{KL}(\cdot)$ is the KL-divergence metric, as the metric to measure the difference between two distributions. The sampled shared embedding is concatenated with image and LiDAR embeddings, along with measurements (velocity, throttle, steer, and brake from the previous frame). This combined information serves as the input for the penalty-based imitation learning module.
III-D Auxiliary Tasks
Auxiliary tasks have been proven to be efficient in many learning approaches. It has two main advantages. 1) By introducing auxiliary tasks, guarantees the important information flows remain in the networks. This information flow is critical for the decision network to make decisions. 2) Auxiliary tasks can guide the direction of gradient descent during training, which can lead the neural network to a more optimal location in the weight space. In this autonomous driving task, we would introduce two extra auxiliary tasks, namely traffic light classification and stop sign classification.
III-D1 Traffic Light Classification
The output of the traffic light decoder should be a vector of 4 which indicates these four states red light, yellow light, green light, and none in the current frame. We then define $y_{\textrm{light}}$ as the ground truth traffic light vector of length 4 and $\hat{y}_{\textrm{light}}$ as the output of the traffic light decoder with the same shape. We use cross-entropy loss $\mathcal{L}_{\textrm{light}}=\mathcal{L}_{\textrm{CE}}(\hat{y}_{\textrm{light}},y_{\textrm{light}})$ for this task.
III-D2 Stop Sign Classification
The output of the stop sign decoder should have a vector of 2 which indicates if a stop sign exists in the current frame. The ground truth stop sign vector of length 2 and the output of the stop sign decoder with the same shape are defined as $y_{\textrm{stop}}$ and $\hat{y}_{\textrm{stop}}$, respectively. Also, cross-entropy loss $\mathcal{L}_{\textrm{stop}}=\mathcal{L}_{\textrm{CE}}(\hat{y}_{\textrm{stop}},y_{\textrm{stop}})$ is used for this task
Note that these two tasks are trained simultaneously with policy generation networks and cross semantics generation tasks. In summary, the final loss function can be calculated as:
$\mathcal{L}_{\textrm{final}}=$ $\mathcal{L}_{\textrm{pl}}+\eta_{1}\mathcal{L}_{\textrm{front}}+\eta_{2}\mathcal{L}_{\textrm{td}}+\eta_{3}\mathcal{L}_{\textrm{light}}+\eta_{4}\mathcal{L}_{\textrm{stop}}+\eta_{5}\mathcal{L}_{\textrm{align}}\text{,}$ (4)
where $\mathcal{L}_{\textrm{pl}}$ is the policy learning loss defined in Equation (1); $\mathcal{L}_{\textrm{front}}$ and $\mathcal{L}_{\textrm{td}}$ are front view and top-down view segmentation losses; $\mathcal{L}_{\textrm{align}}$ is the contractive alignment loss defined in Equation (2); $\mathcal{L}_{\textrm{light}}$ and $\mathcal{L}_{\textrm{stop}}$ are losses for auxiliary tasks, namely traffic light classification and stop sign classification. $\eta_{1}$, $\eta_{2}$, $\eta_{3}$, $\eta_{4}$, $\eta_{5}$ are weights to balance these losses.
III-E Penalty-based Imitation Learning
Our investigation revealed that the objective function used in imitation learning and the metric used for evaluating autonomous driving performance are not consistent with each other. Hence, achieving a low loss in the objective function does not necessarily guarantee a high driving score or successful route completion. We identified two possible factors contributing to this discrepancy.
  • The expert agent still makes mistakes when generating the dataset. Sometimes, the expert agent runs a red light and violates the stop sign rule.
  • The objective function is not sensitive to serious violations of the traffic rules, i.e. the violation of red lights and stop signs. The average objective function loss may not increase too much when violating the traffic rules despite that this violation may cause serious consequences which result in a huge drop in driving score and route completion.
Behavior Cloning (BC) is an imitation learning technique that seeks to replicate the behavior of an expert agent. As a result, the performance of the trained agent is inherently limited by the capabilities of the expert agent. In other words, if the expert agent makes a mistake, the trained agent will learn to replicate that mistake rather than correct it.
Our goal is to modify the objective function used in imitation learning to incorporate traffic rules. Specifically, we plan to introduce additional penalties (higher loss), when the agent generates short-term future waypoints that violate traffic rules during training.
Traffic rules can be expressed as constraint functions that define conditions the optimization problem must meet. In our approach, we focus on three specific aspects of driving behavior: running red lights, ignoring stop signs, and failing to slow down when turning. These are the primary issues we observed with our vanilla imitation learning method. To address these concerns, we propose three corresponding penalties that can be used to quantify and penalize these violations.
III-E1 Red Light Penalty
For the red light violation, we design a red light penalty as follows:
${\mathcal{P}}_{\rm tl}=\mathbb{E}_{(\bm{s},\mathcal{W})\sim\mathcal{D}}[\mathbb{1}_{\rm red}\cdot\sum_{t=1}^{T}c_{i}\cdot{\rm max}\{0,y_{t}-\overline{d}\}]\text{,}$ (5)
where $y_{t}$ represents the $y$ axis value in the $t$-th predicted waypoints $\bm{\hat{w}_{t}}$ generated by the trained agent and $\overline{d}$ denotes the distance between the ego car and the stop line at the intersection along the forward direction of the ego car ($y$ axis). The weight parameter is denoted by $c_{i}$, and the sum of all weight parameters is equal to one. $\mathbb{1}_{\rm red}$ indicates the presence of a red light that may affect the agent in the current frame. $(\bm{s},\mathcal{W})$ represents the combination of observation state $\bm{s}$ and waypoints set $\mathcal{W}$ of one frame, while $\mathcal{D}$ refers to the entire dataset.
When facing red lights, a red light penalty is added based on the distances of the predicted waypoints beyond the stop line at the intersection. If the predicted waypoints fall within the stop line, the penalty remains at zero. Conversely, if the predicted waypoints exceed the stop line, the total distance between those waypoints and the stop line is computed as the red light penalty. The necessary information for calculating the red light penalty, such as traffic light information and stop line location, is pre-processed and stored in each frame of our dataset.
III-E2 Stop Sign Penalty
Similar to the red light penalty, a stop sign penalty is given when the predicted waypoints violate the stop sign rule. The penalty is formalized as follows:
${\mathcal{P}}_{\rm ss}=\mathbb{E}_{(\bm{s},\mathcal{W})\sim\mathcal{D}}[\mathbb{1}_{\rm stopsign}\cdot{\rm max}\{v-\epsilon_{v},0\}]$ (6)
where $v$ is the estimated speed calculated by
$v=\frac{||\bm{\hat{w}_{0}}-\bm{\hat{w}_{1}}||_{2}}{\Delta t}$ (7)
The variables $\bm{\hat{w}_{0}}$ and $\bm{\hat{w}_{1}}$ represent the first and second predicted waypoints, respectively, while $\Delta t$ indicates the time interval between each frame during data collection. The function $\mathbb{1}_{\rm stopsign}$ serves as an indicator for stop sign checking. The maximum speed required to pass stop sign tests is denoted by $\epsilon_{v}$.
An upper-speed limit is established for an area affected by a stop sign, and only speeds lower than this limit are permitted for the agent to pass through. If the agent exceeds this speed limit, a penalty is imposed based on its speed. As the training network only generates predicted waypoints, the velocity estimated from the waypoints is used to compute the stop sign penalty.
III-E3 Speed Penalty
If the agent attempts to turn at excessive speed, a penalty will be enforced. The rationale behind this penalty is based on human driving experience, as it is commonly known that turning at high speeds can increase the risk of collisions with pedestrians or other objects due to reduced braking distance. The speed penalty is defined as follows:
${\mathcal{P}}_{\rm sp}=\mathbb{E}_{(\bm{s},\mathcal{W})\sim\mathcal{D}}[{\rm\sin}(\Delta\delta)\cdot{\max}\{v-v_{\textrm{lb}},0\}]$ (8)
where $\Delta\delta$ denotes the deviation in direction between the current frame and the next frame. As with the stop sign penalty, the desired speed $v$ is defined by Equation 7, and $v_{\rm lb}$ represents the lower speed limit. Any speed below this limit is not subject to the speed penalty.
III-E4 Objctive Function
We trained the network to minimize the difference between the predicted waypoints and the ground truth waypoints, applying an $l_{1}$-norm based loss function. Also, cross semantics generation and auxiliary tasks are trained together with the policy generation task. The final loss function is defined in Equation (4). By applying the penalties, we formalize the constrained optimization:
By applying the penalties, we formalize the constrained optimization:
$\min_{\bm{\Theta}}\quad$ $\mathcal{L}_{\textrm{final}}$ (9)
s.t. $\mathcal{P}_{\rm tl},\mathcal{P}_{\rm ss},\mathcal{P}_{\rm sp}=0$
where $\mathcal{L}_{\textrm{final}}$ is the loss function defined in Equation (4). $\mathbf{\Theta}$ denotes a vector that includes all learnable parameters in the model. It is constrained that the penalties should equal 0 since the predicted waypoints should not violate the traffic rules.
The Lagrange multiplier strategy can be applied here. We introduce three Lagrange Multiplier $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ and the Lagrange function is defined by:
$\min_{\bm{\Theta}}\quad\mathcal{L}_{\textrm{final}}+\lambda_{1}\mathcal{P}_{\rm tl}+\lambda_{2}\mathcal{P}_{\rm ss}+\lambda_{3}\mathcal{P}_{\rm sp}$ (10)
This is the final objective function to optimize. For simplicity, these Lagrange multipliers $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ are considered fixed hyper-parameters. Well-chosen $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ are important for optimization. According to our experiments, too large hyper-parameters influence the behaviors in other scenarios while too small ones are not powerful enough for the agent to obey the corresponding traffic rules. The right part of Figure 1 demonstrates the process of penalty-based imitation learning in detail.
IV Experiments
IV-A Task Description
We focus on a navigation task where the vehicle navigates along predefined routes in various scenarios and areas. The vehicle is guided by GPS signals and situations where GPS signals are low or not available are not being considered. The predefined routes include specific scenarios to test the agent’s ability to respond to emergencies such as obstacle avoidance, other vehicles running red lights, and the sudden appearance of pedestrians on the road. The agent is required to complete the route within a specified time limit, and any time exceeding the limit is considered a failure in terms of route completion.
IV-B Training Dataset
Obtaining realistic driving data is a challenging task. Therefore, we opted to use the Carla simulator [22] to gather training data that has been processed by the expert policy. Our training dataset is the same as TransFuser [6], which comprises approximately 2500 routes through junctions in 8 different towns. These routes have an average length of 100 meters, and around 1000 routes along curved highways with an average length of 400 meters. We utilized the expert agent, similar to TransFuser, to generate this training data. The expert policy we use is the same export policy in TransFuser [6]. For each frame, We collect
  • The forward, left 60 degrees, and right 60 degrees images with the resolutions of $400\times 300$ and their corresponding semantic segmentations.
  • The cloud point of LiDAR with 180 degrees in front of the vehicle and the top-down semantic segmentation.
  • The measurements of the steer, throttle, brake action values, and velocity in the current frame.
  • The traffic light information which includes the stop line position, and traffic light status.
  • The stop sign information indicates if the current frame is influenced by a stop sign.
  • The future waypoints in the next 4 frames.
Note that the interval for each frame is 0.5 seconds, indicating 2FPS collection frequency.
IV-C Metrics
The metrics used to evaluate the behavior of each are Route completion and Infraction Scores.
IV-C1 Route Completion
Route completion (RC) refers to the proportion of the completed route out of the whole route. Suppose $R_{j}$ means the route completion proportion in route $j$ and $N_{\textrm{route}}$ means the number of routes. The RC can be defined by:
$RC=\frac{1}{N_{\textrm{route}}}\sum_{j}^{N_{\textrm{route}}}R_{j}$ (11)
IV-C2 Infraction Score
Infraction Score (IS) is used to measure the driving behavior of the agent. We define $p_{k}$ as the penalty for an infraction instance, $k$ is incurred by the agent and $n_{k}$ is the number of occurrences of infraction instance $k$. Then the Infraction Score can be defined by:
$IS=\prod_{k}^{\{\rm Ped,Veh,Stat,Red,Stop\}}s_{k}^{n_{k}}$ (12)
The infraction instances include collision with a pedestrian, collision with a vehicle, collision with static layout, red light violations, and stop sign violations. The penalty scores $s_{k}$ for them are 0.5, 0.60, 0.65, 0.7, and 0.8, respectively. $n_{k}$ is the number of infractions.
IV-C3 Driving Score
Driving Score (DS) aims to measure the overall driving performance. It is defined by the weighted average of the route completion with an infraction multiplier:
$DS=\frac{1}{N_{\textrm{route}}}\sum_{j}^{N_{\textrm{route}}}R_{j}IS_{j}$ (13)
where $R_{j}$ and $IS_{j}$ are the route completion and infraction score $IS$ (12) for $j$-th test route.
IV-C4 Infractions per km
Infractions per km refer to the average infraction numbers per kilometer. The infractions include collisions with pedestrians, vehicles, and static elements, running a red light, off-road infractions, route deviations, timeouts, and vehicle blocks.
${\rm Infractions\ per\ km}=\frac{\sum_{j}^{N_{\textrm{route}}}\#\ {\rm infractions}_{j}}{\sum_{j}^{N_{\textrm{route}}}l_{j}}$ (14)
where $l_{j}$ is the driven distance (in kilometers) for route $j$.
IV-D Test Results
We use Carla leaderboard - Town05 long benchmarks to evaluate our model. Town05 long benchmark contains 10 routes and all of these routes are over 2.5 km. This benchmark is also used by InterFuser [8] and TransFuser [21]. We also test our model in the Longest6 benchmark which is proposed in TransFuser+ [7].
Baseline. The other baselines we chose to compare with our model are TransFuser+, TransFuser, Geometric Fusion, and LateFusion. TransFuser [21] introduces the Transformer into the multi-sensor fusion architecture to achieve better end-to-end autonomous driving results. TransFuser+ [7], as an extension of TransFuser, leverages several auxiliary losses to ensure important information flows such as traffic light and road line information in the network. InterFuser [8] developed a safety control module to regulate the behaviors of the agent, preventing the agent from violating the traffic rules. LAV [9] utilizes data from both the ego vehicle and surrounding vehicles to enhance the dataset through data augmentation. This is achieved by learning a spatial intermediate representation that remains invariant to different viewpoints.
The evaluation results demonstrated in Table I indicate that our model has a huge increase in driving scores and infraction penalties compared to other SOTA models. As the table demonstrates, our model adheres to traffic rules more effectively in terms of collision, red light violations, and stop sign violations. Table II demonstrates the performance of our model in the longest6 benchmark provided by Transfuser+. Our model continues to surpass all other baselines in driving scores and infraction scores. Additionally, there has been a significant decrease in collisions with pedestrians and vehicles. While our model may not attain the top route completion score, its highest infraction score and driving score suggest that the model embraces a more cautious approach to mitigating traffic rule violations, ultimately leading to the highest driving score.
IV-E Ablation Studies
In this subsection, we analyze the influences of different penalty weights for corresponding traffic rules. As Table III demonstrates, two extra weights for each penalty are selected for comparison. We also provide the results of models without CSG and penalties for comprehensive analysis. We notice that the infractions of traffic lights and stop signs are largely reduced by adding penalties. Our proposed multi-sensor fusion technology (CSG) also decreases the possibility of hitting obstacles such as vehicles, pedestrians, and other statics. The results of different penalty weights are also listed in the table for comparison. We found that assigning greater weight to more severe violations will increase the performance of our model. For instance, we apply greater penalties for the red light and the stop sign violations compared to overspeeding by turning, since those two violations cause more serious consequences.
Model Driving score Route compl. Infrac. score Collision pedestrian Collision vehicle Collision static Red light infraction Stop Sign Infraction
[…]
P-CSG(Ours) $\textbf{59.10}\pm 4.07$ $86.84\pm 2.99$ $\textbf{0.68}\pm 0.06$ $0.01\pm 0.01$ $\textbf{0.07}\pm 0.05$ $0\textbf{0.00}\pm 0$ 0.00 $0.04\pm 0.02$ $0\textbf{0.00}\pm 0$ 0.00
TransFuser [21] $34.50\pm 2.54$ $61.16\pm 4.75$ $0.56\pm 0.06$ $0.01\pm 0.01$ $0.58\pm 0.07$ $0.38\pm 0.05$ $0.12\pm 0.03$ $0.05\pm 0.02$
TransFuser+ [7] $36.19\pm 0.90$ $70.13\pm 6.80$ $0.51\pm 0.03$ $0\textbf{0.00}\pm 0$ 0.00 $0.40\pm 0.13$ $0.04\pm 0.03$ $0.11\pm 0.10$ $0.04\pm 0.01$
InterFuser [8] $50.64\pm 3.51$ $89.13\pm 4.12$ $0.57\pm 0.05$ $0\textbf{0.00}\pm 0$ 0.00 $0.09\pm 0.04$ $0.01\pm 0.01$ $\textbf{0.02}\pm 0.01$ $0.04\pm 0.01$
LAV [9] $45.20\pm 6.36$ $\textbf{91.55}\pm 5.61$ $0.49\pm 0.06$ $0\textbf{0.00}\pm 0$ 0.00 $0.92\pm 0.42$ $0.33\pm 0.50$ $0.28\pm 0.28$
TABLE I: Carla Leaderboard - Town 05 Long Benchmark. In our assessment of the Town05 long benchmark, we analyze several performance indicators including the Driving Score, Route Complication, Infraction Score, instances of collisions, violations of red light signals, and violations related to stop signs. These metrics are precisely outlined by the Carla leaderboard [22]. It’s important to highlight that all alternate baseline models undergo retraining and testing within our dedicated CARLA environment. Each model’s evaluation is conducted thrice, employing distinct random seeds on every occasion. The performance data of LAV is collected from their paper.
Model Driving score Route compl. Infrac. score Collision pedestrian Collision vehicle Collision static Red light infraction Stop Sign Infraction
[…]
P-CSG(Ours) $\textbf{49.21}\pm 2.93$ $86.24\pm 2.46$ $\textbf{0.56}\pm 0.04$ 0.00 0.18 0.24 0.21 0.00
WOR [27] $20.53\pm 3.12$ $48.47\pm 3.86$ $0.56\pm 0.03$ 0.18 1.05 0.37 1.28 -
LAV [9] $32.74\pm 1.45$ $70.36\pm 3.14$ $0.51\pm 0.02$ 0.16 0.83 0.15 0.96 -
Late Fusion $22.47\pm 3.71$ $83.30\pm 3.04$ $0.27\pm 0.04$ 0.05 4.63 0.28 0.11 -
TransFuser+ [7] $47.30\pm 5.72$ $\textbf{93.38}\pm 1.20$ $0.50\pm 0.06$ 0.03 2.45 0.07 0.16 -
TABLE II: Carla Leaderboard - Longest6 Benchmark We assess our model with Longest6 benchmark provided by Transfuser+ [7]. This benchmark is also under the Carla leaderboard framework. The metrics of the driving score, route complication, infraction score, collisions, violations of red light signals, and violations related to stop signs are the same as the metrics in the Carla leaderboard. The performance data of all other baselines is collected in their paper.
Model Driving score Route compl. Infrac. score Collision pedestrian Collision vehicle Collision static Red light infraction Stop Sign Infraction
[…]
P-CSG(Ours) $\textbf{59.10}\pm 4.07$ $86.84\pm 2.99$ $\textbf{0.68}\pm 0.06$ $0.01\pm 0.01$ $\textbf{0.07}\pm 0.05$ $0\textbf{0.00}\pm 0$ 0.00 $0.09\pm 0.02$ $0\textbf{0.00}\pm 0$ 0.00
No CSG $45.67\pm 5.51$ $84.82\pm 2.25$ $0.52\pm 0.10$ $0.02\pm 0.01$ $0.14\pm 0.03$ $0.03\pm 0.02$ $0.09\pm 0.02$ $0.01\pm 0.01$
No Penalty $37.32\pm 3.58$ $78.41\pm 11.32$ $0.54\pm 0.21$ $0.02\pm 0.01$ $0.11\pm 0.03$ $0.02\pm 0.01$ $0.11\pm 0.03$ $0.04\pm 0.02$
$\lambda_{1}=0.7$ $54.36\pm 1.78$ $84.37\pm 5.01$ $0.64\pm 0.04$ $0.01\pm 0.01$ $0.18\pm 0.02$ $0\textbf{0.00}\pm 0$ 0.00 $\textbf{0.08}\pm 0.01$ $0.01\pm 0.00$
$\lambda_{1}=0.3$ $56.80\pm 2.22$ $89.60\pm 1.90$ $0.60\pm 0.02$ $0.01\pm 0.01$ $0.18\pm 0.01$ $0\textbf{0.00}\pm 0$ 0.00 $0.13\pm 0.01$ $0.01\pm 0.01$
$\lambda_{2}=0.07$ $51.54\pm 1.50$ $81.49\pm 1.42$ $0.63\pm 0.02$ $0.02\pm 0.01$ $0.13\pm 0.10$ $0\textbf{0.00}\pm 0$ 0.00 $0.13\pm 0.01$ $0.01\pm 0.01$
$\lambda_{2}=0.03$ $54.59\pm 2.88$ $88.12\pm 1.95$ $0.64\pm 0.02$ $0.01\pm 0.01$ $0.11\pm 0.01$ $0\textbf{0.00}\pm 0$ 0.00 $\textbf{0.08}\pm 0.06$ $0.01\pm 0.01$
$\lambda_{3}=0.7$ $55.98\pm 3.00$ $\textbf{89.81}\pm 3.33$ $0.63\pm 0.02$ $0.01\pm 0.01$ $0.09\pm 0.02$ $0\textbf{0.00}\pm 0$ 0.00 $0.09\pm 0.04$ $0\textbf{0.00}\pm 0$ 0.00
$\lambda_{3}=0.3$ $51.98\pm 1.93$ $84.66\pm 3.53$ $0.63\pm 0.01$ $0\textbf{0.00}\pm 0$ 0.00 $0.08\pm 0.03$ $0.02\pm 0.00$ $0.10\pm 0.02$ $0.01\pm 0.01$
TABLE III: Ablation Study - Town 05 Long Benchmark The ablation study focuses on significant hyper-parameters, namely $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$, which correspond to the penalty weights associated with the red light, speed, and stop sign, respectively. Each model is evaluated three times with random seeds. Our optimal model employs default weights of $\lambda_{1}=0.5$, $\lambda_{2}=0.05$, and $\lambda_{3}=0.5$. During the ablation study, we isolate the impact of a single hyper-parameter by modifying it while keeping the others constant.
V Robustness Study
Attack Model Driving score Route compl. Infrac. score Collision pedestrian Collision vehicle Collision static Red light infraction Stop Sign Infraction
[…]
FGSM Attack $(\epsilon=0.01)$ P-CSG(Ours) $\mathbf{12.80}\pm 1.35$ $21.70\pm 2.87$ $\mathbf{0.59}\pm 0.02$ $\mathbf{0.00}\pm 0.00$ $\mathbf{0.22}\pm 0.23$ $2.40\pm 1.00$ $\mathbf{0.07}\pm 0.11$ $\mathbf{0.00}\pm 0.00$
Transfuser+ $2.69\pm 0.73$ $42.44\pm 3.65$ $0.22\pm 0.06$ $\mathbf{0.00}\pm 0.00$ $1.23\pm 0.31$ $0.26\pm 0.17$ $0.08\pm 0.03$ $0.04\pm 0.02$
Interfuser* $8.37\pm 1.06$ $\mathbf{72.04}\pm 4.15$ $0.15\pm 0.04$ $0.00\pm 0.01$ $0.64\pm 0.27$ $\mathbf{0.10}\pm 0.03$ $0.22\pm 0.03$ $0.07\pm 0.01$
Dot Attack P-CSG(Ours) $\mathbf{36.60}\pm 3.17$ $\mathbf{75.53}\pm 6.43$ $\mathbf{0.56}\pm 0.07$ $0.01\pm 0.02$ $\mathbf{0.06}\pm 0.03$ $\mathbf{0.00}\pm 0.00$ $\mathbf{0.10}\pm 0.04$ $\mathbf{0.01}\pm 0.00$
Transfuser+ $0.53\pm 0.21$ $54.06\pm 5.48$ $0.02\pm 0.01$ $\mathbf{0.00}\pm 0.00$ $2.53\pm 0.72$ $0.99\pm 0.16$ $0.21\pm 0.04$ $0.04\pm 0.01$
Interfuser* $12.56\pm 2.15$ $73.62\pm 7.32$ $0.19\pm 0.02$ $0.01\pm 0.01$ $0.68\pm 0.26$ $\mathbf{0.00}\pm 0.00$ $0.21\pm 0.02$ $0.09\pm 0.02$
TABLE IV: Attack Results. This table presents a comparative analysis of the performance and robustness of autonomous driving models under two types of adversarial attacks: FGSM [28] and Dot Attacks [29]. Driving Score serves as the primary metric of overall performance, while the infraction score aggregates various safety-related factors. P-CSG, our proposed model, consistently outperforms the other models across these metrics. Interfuser* has the same PID Controller as P-CSG and Transfuser+ to ensure a fair comparison.
Given the life-critical nature of autonomous driving, it becomes essential to examine the robustness of neural network models against such vulnerabilities.
In this regard, we conduct a comprehensive comparative analysis to evaluate the robustness of our model. We benchmark its performance against two state-of-the-art architectures—Transfuser+ [7] and Interfuser [8]. We focus on two specific types of white-box sensor attacks. These are Fast Gradient Sign Method (FGSM) Attack [28] and Dot Attack [29].
FGSM Attack [28]: FGSM Attack is notable for their computational efficiency and minimal perceptibility to the human eye. This approach exploits the gradient of the model’s loss function with respect to the input data.
$\bm{\eta}=\epsilon_{\textrm{p}}\operatorname{sign}(\nabla_{I}J(\bm{\Theta},\bm{I},y)).$ (15)
$\bm{\tilde{I}}=\bm{I}+\bm{\eta}$ (16)
where $\bm{I}$ is the original image input, and $y$ is the ground true label. The loss function is $J(\bm{\Theta},\bm{I},y)$, and its gradient with respect to ${\bm{I}}$ is $\nabla_{\bm{I}}J(\bm{\Theta},\bm{I},y)$. The hyperparameter $\epsilon_{\textrm{p}}$ controls perturbation magnitude, leading to adversarial input $\bm{\tilde{{I}}}$. $\operatorname{sign}(\cdot)$ is the sign indicator which has the value of 1 for positive value and -1 for negative value. The method creates subtle but impactful perturbations in autonomous driving.
Figure 2: Qualitative Attack Results on P-CSG. (a) Original RGB input, (b) Dot Attack with nine trained dots, (c) FGSM Attack with $\epsilon=0.01$. The subtle FGSM perturbation in (c) is hard to spot compared to (a), showcasing the method’s ability to create changes that are invisible to the human eye but still mislead models.
Dot Attack[29] : Dot Attack targets deep learning algorithms by introducing minor, physical disruptions to camera systems. Specifically, a sticker with a designed dot pattern is placed on the camera lens.
Given an RGB image input $\bm{I}$, semi-transparent dots are strategically placed. We first consider the perturbation function for a single dot on the image $\bm{I}$. The perturbed image $\bm{\tilde{I}}$ at pixel location $(x_{i},x_{j})$ can be written as
$\begin{split}\bm{\tilde{I}}[x_{i},x_{j}]&=f_{0}(\bm{\phi_{0}};x_{i},x_{j})\\ &=(1-h(x_{i},x_{j}))\cdot\bm{I}(x_{i},x_{j})+h(x_{i},x_{j})\cdot\gamma\text{,}\end{split}$ (17)
where
$h(x_{i},x_{j})$ $=\alpha\cdot\exp\left(-\left(\frac{(x_{i}-x_{i}^{(c)})^{2}+(x_{j}-x_{j}^{(c)})^{2}}{r^{2}}\right)^{\beta}\right)$ (18)
$\bm{\phi_{0}}$ $=(\gamma,(x_{i}^{(c)},x_{j}^{(c)}),r,\alpha,\beta)\text{.}$ (19)
Here, $h(x_{i},x_{j})$ quantifies the balance between the original color at pixel $\bm{I}[x_{i},x_{j}]$ and a learnable color $\gamma$. This balance reduces with increasing distance from the center of the dot $\left(x_{i}^{(c)},x_{j}^{(c)}\right)$. $\bm{\phi_{0}}$ denotes the learnable parameters.
By composing $N_{\textrm{dot}}$ of these single-dot perturbations:
$f(\bm{\phi_{0}},\bm{\phi_{1}},...,\bm{\phi_{N_{\textrm{dot}}}};x_{i},x_{j})=f(\bm{\phi_{0}};x_{i},x_{j})\circ...\circ f(\bm{\phi_{N_{\textrm{dot}}}};x_{i},x_{j})\text{,}$ (20)
the pixel at the location $(x_{i},x_{j})$ of the final perturbed image $\bm{\tilde{I}}$ can be formalized as:
$\bm{\tilde{I}}[x_{i},x_{j}]=f_{0}(\bm{\phi_{0}},\bm{\phi_{1}},...,\bm{\phi_{N_{\textrm{dot}}}};x_{i},x_{j})\text{.}$ (21)
These parameters are learned by increasing the loss of the model, leading to poor performance of the trained model. In Fig. 2(b), the resulting dots mimic real-world lens stains, displaying a color gradient. Unlike computational attacks like FGSM, Dot Attacks are easily deployed in the real world, requiring only the application of a sticker to the camera lens.
V-A Experimental Design
In our experiments conducted within the CARLA simulation environment, we utilized a predefined trajectory, termed Town05 Long. To ensure a fair comparison across three distinct deep learning models—Transfuser+, Interfuser, and P-CSG—a consistent PID Controller was applied to all models. This uniform approach was designed to specifically assess the inherent robustness of each model under different adversarial attacks, by eliminating variations in control mechanisms.
FGSM Attacks: Our goal is to prompt immediate deviations in the vehicle’s trajectory by assigning target waypoints for the next four timesteps at $(-1.0,-2.0)$, $(-2.0,-4.0)$, $(-3.0,-6.0)$, and $(-4.0,-8.0)$, relative to the vehicle’s current position. These are used to formulate a dedicated loss function, from which adversarial perturbations are generated and applied to the input RGB images. We set the perturbation magnitude $\epsilon$ to 0.01 for these experiments.
Dot Attacks: Unlike FGSM Attacks, which necessitate per-frame computations, Dot Attacks employ a pre-trained, universally applicable Dot Mask. Our experimental procedure initiated with the nine-dots mask training phase, followed by its application within the CARLA simulation framework. This structured approach permitted us to evaluate the practicality and effectiveness of Dot Attack under conditions that closely mimic real-world scenarios.
V-B Results
In our evaluation, the Infraction Score serves as the primary metric, encapsulating various collision scenarios and traffic violations to comprehensively assess safety. In the context of adversarial attacks, a higher infraction score is crucial for the safety of both passengers and pedestrians. As shown in Table IV, our model, P-CSG, demonstrates superior performance in two distinct adversarial scenarios: high-intensity, per-frame FGSM Attack, and uniformly applied, low-intensity Dot Attack.
Specifically, P-CSG achieved an Infraction Score that was at least $2.5$ times higher than that of competing models in both attack scenarios tested. Under the low-intensity Dot Attack, P-CSG not only maintained high safety standards, indicated by a higher Infraction Score but also completed more routes, as reflected in a higher Route Completion metric. Conversely, in scenarios involving high-intensity FGSM Attacks, P-CSG implemented a more conservative strategy aimed at enhancing security. This is shown by its higher Infraction Score, even though it led to fewer completed routes.
Most importantly, P-CSG consistently reaches the highest cumulative Driving Score in both types of attacks, confirming its robustness and adaptability in different challenging conditions.
VI Conclusion
In this paper, we improve the fusion technologies and policy learning approaches based on penalties for autonomous driving. Additionally, we compare our proposed model with other baselines under Dot Attacks and FGSM Attacks, demonstrating outstanding performance and robustness. We aspire for our proposed penalty-based imitation learning approach to introduce a fresh perspective into the domain of end-to-end autonomous driving, aimed at enhancing the compliance of autonomous agents with traffic rules. Furthermore, additional research is necessary to enhance the robustness of autonomous driving systems, particularly in countering attacks such as Dot Attacks.

### v3

@@
-What Matters to Enhance Traffic Rule Compliance of Imitation Learning for Automated Driving
+
 Abstract
-More research attention has recently been given to end-to-end autonomous driving technologies where the entire driving pipeline is replaced with a single neural network because of its simpler structure and faster inference time.
-Despite this appealing approach largely reducing the components in the driving pipeline, its simplicity also leads to interpretability problems and safety issues [1].
-The trained policy is not always compliant with the traffic rules and it is also hard to discover the reason for the misbehavior because of the lack of intermediate outputs.
-Meanwhile, sensors are also critical to autonomous driving’s security and feasibility to perceive the surrounding environment under complex driving scenarios.
-In this paper, we proposed P-CSG, a penalty-based imitation learning approach with cross semantics generation sensor fusion technologies to increase the overall performance of end-to-end autonomous driving.
+End-to-end autonomous driving, where the entire driving pipeline is replaced with a single neural network, has recently gained research attention because of its simpler structure and faster inference time.
+Despite this appealing approach largely reducing the complexity in the driving pipeline, it also leads to safety issues because the trained policy is not always compliant with the traffic rules.
+In this paper, we proposed P-CSG, a penalty-based imitation learning approach with contrastive-based cross semantics generation sensor fusion technologies to increase the overall performance of end-to-end autonomous driving.
 In this method, we introduce three penalties - red light, stop sign, and curvature speed penalty to make the agent more sensitive to traffic rules.
-The proposed cross semantics generation helps to align the shared information from different input modalities.
-We assessed our model’s performance using the CARLA leaderboard - Town 05 Long benchmark and Longest6 Benchmark, achieving an impressive driving score improvement.
-Furthermore, we conducted robustness evaluations against adversarial attacks like FGSM and Dot attacks, revealing a substantial increase in robustness compared to baseline models.
-More detailed information, such as code base resources, and videos can be found at https://hk-zh.github.io/p-csg-plus.
-Index Terms:
-Imitation Learning, Multi-modality Sensor Fusion, End-to-end Autonomous Driving
-I Introduction
-End-to-end autonomous driving integrates the perception and decision layers into one deep neural network.
-Enhancing the policy learning algorithm and improving perception abilities are two crucial aspects that can significantly enhance the overall performance of autonomous driving.
-Imitation learning [2] and reinforcement learning [3] are two main learning paradigms unutilized end-to-end autonomous driving systems for the decision-making part.
-Although both approaches have been successful in the field of autonomous driving, they each have their own drawbacks.
-Imitation learning only learns the behaviors limited to the training data.
-This means that the model may struggle to generalize to new and unfamiliar situations.
-The quality of the expert policy also has a significant impact on the performance of imitation learning.
-The design of the reward function plays a critical role in the effectiveness of reinforcement learning for autonomous driving systems, and creating an appropriate reward function can be a complex task.
-Additionally, there is a risk that reinforcement learning models may learn unexpected and potentially dangerous behaviors in such safety-critical systems.
-In this article, we propose a penalty-based imitation learning paradigm that introduces the "penalty"/"reward" concept from reinforcement learing into imitation learning and aims to 1) achieve higher learning efficiency compared to reinforcement learning (learning from the expert demonstration) and 2) make the agent more sensitive to the traffic rules by involving some penalties designed in a simple manner.
-The perception component in the neural network is used to extract essential information about the surrounding environment.
-Despite some Lidar-based approaches [4] [5], which leverages Lidar sensor input and HD maps, demonstrate impressive performance, relying on HD maps is not a feasible option as it requires substantial resources to create and maintain.
+The proposed cross semantics generation helps to align the shared information of different input modalities.
+We assessed our model’s performance using the CARLA Leaderboard - Town 05 Long Benchmark and Longest6 Benchmark, achieving 8.5% and 2.0% driving score improvement compared to the baselines.
+Furthermore, we conducted robustness evaluations against adversarial attacks like FGSM and Dot attacks, revealing a substantial increase in robustness compared to other baseline models.
+Keywords:
+Imitation Learning Multi-modality Sensor Fusion End-to-end Autonomous Driving
+1 Introduction
+End-to-end autonomous driving [20] integrates the perception and decision-making layers into one deep neural network.
+The perception component extracts essential information about the surrounding environment.
+Despite some approaches [4] [32], which leverage LiDAR sensor input and HD maps and demonstrate impressive performance, relying on HD maps is not a feasible option as it requires substantial resources to create and maintain.
 They may not be universally available for all areas and regions.
-For this reason, an increasing number of research [6] [7] [8] [9] recently concentrate on multi-modality sensors (e.g Lidar and Camera) based approaches.
-Sensor fusion technologies are widely used in 2D object and 3D object detection.
-Classified by the types of point cloud representation of Lidar input, there exist three main approaches, namely point-based [10] [11], voxel-based [12] [13] [14] [15], and range-view-based [16] [17] [18] [19] approaches.
-More recently, with the popularity of attention mechanisms, many researchers are trying to use Transformer [20] to integrate multimodal information [7] [21] [8].In this paper, We shift our approach by extracting and aligning shared information from diverse modalities, enabling the model to extract the global context within multi-modalities
-We propose Penalty-based Imitation Learning with Cross Semantics Generation (P-CSG).
-We evaluate our model on Carla Leaderboard - Town 05 Long Benchmark and Longest6 Benchmark.
-It achieves state-of-the-art performance, especially in collision and traffic rule violations.
-We also analyze the performance and robustness of our autonomous driving model under two types of adversarial attacks: FGSM and Dot attacks.
-The results show that our model outperforms the other baselines from the perspective of safety.
+Recent research has concentrated on multi-modality sensor techniques utilizing both LiDAR [24, 15] and cameras [35, 12, 33, 6].
+To manage the complexity of LiDAR input, various approaches have been developed, including point-based [44], voxel-based [34, 46, 48, 41], and range-view-based approaches [14, 8, 36].
+More recently, with the popularity of attention mechanisms, many researchers are trying to use Transformer [39, 43] to integrate multimodal information [12] [29] [33].
+Despite its potential, the Transformer architecture on a large scale leads to a significant increase in both the training and inference time.
+In this paper, we shift our approach by extracting and aligning shared information from diverse modalities by contrastive learning, enabling the model to extract the global context within multi-modalities.
+Refer to caption(a)(b)
+Figure 1: (a) Common imitation learning approach could learn the wrong behavior which violates the traffic rule.
+(b) Our approach penalizes the predicted waypoints violating traffic rules to enhance the agent’s traffic rule adherence.
+Imitation learning [1] and reinforcement learning [37] are two main learning paradigms utilized by end-to-end autonomous driving systems for the decision-making part.
+Reinforcement learning suffers from the low sample efficiency and the complexity of reward signs.
+It may also learn risky actions, leading to accidents and unsafe driving.
+Due to these shortages, other researchers have turned to imitation learning approaches [39, 12, 29, 33].
+However, these approaches still lack effective mechanisms to ensure the agent to comply the traffic rules.
+For one thing, the expert policy to generate demonstrations may still make mistakes when collecting the training data.
+Also, the metrics for autonomous driving (e.g., traffic rule violations) and the objectives of imitation learning (minimize the difference between the predicted and ground truth waypoints) remain disparate, indicating that a low loss in the learning objective does not guarantee optimal performance of the agent in the testing environments.
+To overcome the above-mentioned limitations, we propose Penalty-based Imitation Learning with contrastive-based Cross Semantics Generation (P-CSG).
+Our approaches penalize the behavior (higher loss) that violates the traffic rules during the imitation learning process to make the agent better adhere to these traffic rules (cf.
+Figure 1).
+Meanwhile, we adopt a contrastive-based sensor fusion technology that aligns the shared information of multi-modalities so that the model can better understand the global context within multi-modalities.
+We evaluated our model on the CARLA Leaderboard - Town 05 Long Benchmark and Longest6 Benchmark.
+It achieves outstanding performance, especially in collision and traffic rule violations.
+We also analyze the performance and robustness of our autonomous driving model under two types of adversarial attacks: Fast Gradient Sign Method (FGSM) Attack [17] and Dot attacks [22].
+The results show that our model outperforms the other baselines from the safety perspective.
 Our main contributions to this paper can be summarized as follows:
-  • Our innovative multi-sensor fusion technique involves aligning shared information from various modalities, thus simplifying the process of extracting the global context across diverse modalities.
+  • Our innovative multi-sensor fusion technique involves aligning shared information from various modalities, thus extracting the global context across diverse modalities.
   • We proposed a penalty-based imitation learning approach that leverages constraint optimizations to make the imitation learning model more sensitive to traffic rule violations.
-II Related Works
-There are currently two primary methods used in autonomous driving technology: modular and end-to-end approaches.
-The modular approach involves a detailed series of software modules that work together to control the vehicle.
-By employing a modular approach, engineering teams can focus on specific and well-defined subtasks and make enhancements independently throughout the entire system.
-This allows the system to remain operational as long as the intermediate outputs remain functional.
-On the other hand, the end-to-end approach considers the entire driving process as one comprehensive learning task.
-It leverages the ability of the neural network to learn task-specific features and perform task-specific actions.
-Contrary to the modular approach, the agent learns without relying on explicit rules or hand-engineered features, namely the human-defined information bottlenecks.
-This means that the vehicle can adapt to different driving scenarios and environments and potentially achieve better performance than systems that rely on traditional rule-based approaches.
-End-to-end approaches have shown great success in computer vision, such as object detection, object tracking, and semantic segmentation.
-The success of these tasks builds a solid foundation for end-to-end autonomous driving.
-It is reasonable to assume that in the future, the end-to-end approach will also have the ability to operate a vehicle autonomously.
-Lately, there has been a growing trend in research to utilize behavior cloning for training agents to accomplish autonomous driving tasks.
-This involves the agent learning from demonstrations provided by experts in the field [7] [8] [6] [9].
-They collect the demonstration in CARLA [22], a simulation urban environment in which the researchers and developers can test and evaluate their algorithms and models for autonomous driving.
-They employ strategies like enhancing multi-modality sensor fusion [7] [8], introducing safety module [8] to avoid dangerous behavior and improve the overall performance of autonomous driving.
-In our approach, penalties for traffic rules are introduced to ensure the safety of driving.
-We also introduce the cross semantics generation, which generates the semantic information of the other modality, to capture the global context across different modalities.
-III Methodologies
+  • We evaluate our method’s performance under adversarial attacks w.r.t. FSGM and Dot Attacks. The results show that our model adopts a more cautious strategy and has a more robust performance.
+The rest of this paper is organized as follows: Section 2 offers a detailed review of related literature.
+We present our methodology in Section 3.
+Section 4 provides the experiment results and discussions.
+Section 5 conduct robustness study against adversarial attacks.
+We conclude our paper in Section 6.
+2 Related Works
+Imitation learning [35, 29, 6, 12, 33, 2] are widely used in the field of autonomous driving [9, 10, 11, 38].
+Many approaches use the CARLA simulator [13] to collect training data and test the performance of the trained agent in the simulation environment.
+To improve the agent’s overall performance, many approaches investigate the multi-sensor fusion technologies and safety mechanisms.
+2.1 Multi-sensor Fusion Technologies
+Multi-sensor fusion has received much research attention in the field of end-to-end autonomous driving.
+Recent works [42, 3, 47, 28], complementing camera images with depth and semantics has shown the potential to improve the overall driving performance.
+More research attention has been given to LiDAR and camera input fusion since they are complementary to each other in terms of representing the scene and are readily available in autonomous driving systems.
+Early works such as LateFusion [35] concatenate the feature embeddings from different modalities and use a Multi-Layer-Perception (MLP) network to weigh, select, and fuse the concatenated features.
+TransFuser [29] leverage Transformer-based approaches to achieve multi-modal global context.
+They apply cross-attention to the feature embeddings from different modalities in different layers.
+TransFuser+ [12], as an extension of TransFuser, introduces auxiliary tasks: depth prediction and semantic segmentation are performed with image information, while HD map prediction and vehicle object detection are handled with bird’s eye view (BEV) LiDAR information.
+These auxiliary tasks serve to illuminate the inner workings of the entire network.
+Similarly, InterFuser [33] also adopts the Transformer-based sensor fusion approach.
+Contrastive Learning, which is widely used in vision language models (VLMs) [30, 23], demonstrates its strong performance in multi-modal information handling.
+Some recent works [26, 45, 25] focus on using contrastive learning for multi-modal information fusion.
+In this paper, we investigate contrastive learning for LiDAR and camera input fusion and propose a contrastive-based cross semantic generation sensor fusion method to extract the global context within the modalities of LiDAR and camera.
+2.2 Safety Mechanism
+To bolster the safety of autonomous driving agents, researchers in [12] concentrate on refining the expertise of the expert agent.
+By filtering out incorrectly demonstrated actions executed by the expert policy, the agent avoids being misled and can more effectively acquire skills from the demonstrations.
+LAV [6] supervises waypoint outputs with additional data by making predictions for other nearby agents.
+InterFuser [33] developed a safety control module to regulate the agent’s behaviors, preventing the agent from violating the traffic rules.
+However, this control module heavily depends on hand-designed heuristics and can vary in different systems.
+In our paper, we introduce the ‘penalty’ concept to the IL framework, which incentivizes the trained agent to adopt safer driving behaviors in an end-to-end manner.
+2.3 Combining IL and RL
+Methods such as DQfD[19], DDPGfD[40], and DAPG[31] have shown that integrating IL with RL can help address exploration challenges in domains with sparse rewards.
+Other offline RL approaches, such as TD3+BC [16] and CQL [21], combine RL and IL objectives to regularize Q-learning updates and prevent overestimation of out-of-distribution values.
+In the autonomous driving field, Lu et al.
+[27] proposed behavior cloned soft actor-critic (BC-SAC), which incorporates a behavioral cloning objective into the SAC objective.
+Our approach starts with an IL perspective by directly incorporating the concept of penalties and rewards from RL into IL.
+We model penalties as constraints and violating these constraints leads to higher loss.
+Q-function learning is not included in our approach.
+3 Methodologies
 In this section, we propose a multi-sensor fusion approach and a penalty-based imitation learning method for end-to-end autonomous driving.
-Feature ExtractionCross Semantic GenerationAuxiliaryPenalty-based Imitation LearningRefer to captionRefer to captionRefer to captionptResnet50ptResnet18pt512pt512ptDecoderRefer to captionRefer to captionptMLPptMLPContractive Losspt128Refer to captionRefer to captionRefer to captionpt4Measurementspt1156Refer to captionRed Light PenaltyRefer to captionStop Sign PenaltyRefer to captionSpeed PenaltyptMLPpt64ptGRUptptGoalLocationptptptWaypointsptPIDRefer to captionRefer to captionPenalties
-Figure 1: This is an overview of our proposed Penalty-based Imitation Learning with Cross Semantics Generation.
-Let’s start with the first feature extraction part.
-The top-down LiDAR pseudo image is fed into a ResNet18 architecture to extract the features from the LiDAR input.
-The front camera image is fed into a ResNet50 architecture to extract the features from the camera input.
-Then, the Cross Semantic Generation is conducted to cross-generate the front and top-down semantic segmentations in the frame of the front image and LiDAR pseudo image.
-We apply the contrastive loss the align the shared features of the camera and LiDAR inputs.
-The resampled shared embedding from latent space will be used to generate the waypoints.
-The features extracted from the front camera image are also leveraged to perform auxiliary tasks ’stop sign indication’ and ’red light indication’.
-In terms of penalty-based imitation learning, the LiDAR features, camera features, shared features, and measurements are concatenated and fed into an MLP network to reduce the dimension.
-A GRU structure is followed to recurrently generate the waypoints, also taking the goal location as input.
-The generated waypoints will be applied with the three penalties, namely, red light penalty, stop sign penalty, and speed penalty.
-The task of the PID controller is to map the waypoints into the action of steer, throttle, and brake.
-III-A Problem Setting
-The task we concentrate on is point-to-point navigation in an urban setting where the goal is to complete a route with safe reactions to dynamic agents such as moving vehicles and pedestrians.
-The traffic rules like red lights and stop signs should be followed.
-We consider the problem of learning a goal-conditioned policy $\pi(\bm{a}|\bm{s},\bm{g})$ that outputs action $\bm{a}\in\mathcal{A}$, conditioned on the current observed state $\bm{s}\in\mathcal{S}$ and a high-level goal location $\bm{g}$ provided by GPS.
+3.1 Problem Setting
+We concentrate on point-to-point navigation in an urban setting where the goal is to complete a route with safe reactions to dynamic agents such as moving vehicles and pedestrians.
+Traffic rules should be followed, such as red lights and stop signs.
+We consider the problem of learning a goal-conditioned policy $\pi(\bm{a}|\bm{s},\bm{g})$ that outputs action $\bm{a}\in\mathcal{A}$, conditioned on the current observed state $\bm{s}\in\mathcal{S}$ and a goal location $\bm{g}$ provided by GPS.
 The environment we used to train and test the performance of the trained agent is CARLA.
 The specific environment setting we used can be characterized by the following statements:
-  • A state space $\mathcal{S}\subset\{(\mathbb{R}^{H_{\textrm{c}}\times W_{\textrm{c}}\times 3},\mathbb{R}^{H_{\textrm{L}}\times W_{\textrm{L}}\times 2},\mathbb{R}^{4})\}$ which is the combination of the observation of camera with the shape of $H_{\textrm{c}}\times W_{\textrm{c}}\times 3$, bird’s eye view LiDAR input with the shape of $H_{\textrm{L}}\times W_{\textrm{L}}\times 2$, and measurements of current throttle, brake, steer, and velocity.
-  • A multidimensional action space $\mathcal{A}\subset\mathbb{R}^{3}$. The action space contains the parameters of steer, throttle, and brake to drive the agent to finish tasks.
+  • A state space $\mathcal{S}\subset\{(\mathbb{R}^{H_{\textrm{c}}\times W_{\textrm{c}}\times 3},\mathbb{R}^{H_{\textrm{L}}\times W_{\textrm{L}}\times 2},\mathbb{R}^{4})\}$ which is the combination of the observation of camera with the shape of $H_{\textrm{c}}\times W_{\textrm{c}}\times 3$, bird’s eye view (BEV) LiDAR pseudo image input with the shape of $H_{\textrm{L}}\times W_{\textrm{L}}\times 2$, and measurements of current throttle, brake, steer, and speed.
+  • A multidimensional action space $\mathcal{A}\subset\mathbb{R}^{3}$.
+The action space contains the parameters of steer, throttle, and brake to drive the agent to finish tasks.
+Instead of directly predicting the next action of throttle, steer, and brake, we estimate the future waypoints $\mathcal{W}$ of the ego-vehicle in BEV space, which is centered at the ego vehicle’s current coordinate frame.
+Then, we use two fine-tuned PID controllers for lateral and longitudinal control to obtain steer, throttle, and brake values from the predicted waypoints.
   • A goal space consists of high-level goal location provided as GPS coordinates.
-III-A1 Observation Space
-Similar to prior works [23], [24] [7] [21], the LiDAR point cloud data is transformed into a 2-bin histogram by mapping it onto a 2D BEV grid of fixed resolution.
-To encompass a 32m x 32m BEV grid, we limit our analysis to points located within 32m in front of the ego vehicle and 16m on each side.
-We partition the grid into blocks of 0.125m $\times$ 0.125m, which gives us a resolution of 256 $\times$ 256 pixels.
-The height dimension is discretized into two bins, distinguishing points that are either on/below or above the ground plane.
-This results in a two-channel pseudo image input $\mathbb{R}^{256\times 256\times 2}$.
-In terms of RGB input, we leverage three cameras positioned to face forward, 60 degrees to the left, and 60 degrees to the right.
-These cameras have a horizontal field of view (FOV) of 120 degrees.
-The image from the cameras has a resolution of $400\times 300$ pixels.
-The forward image is cropped to $300\times 160$ pixels and the 60 degrees to the left and right images are cropped to $160\times 234$ pixels.
-we want to use cropping to eliminate radial distortion at the edges.
-We combine these three cropped images into a single image input with a resolution of $768\times 160$ pixels.
-In addition, the current throttle, brake, steering, and velocity measurements are included in the observation space to provide a point of reference for the agent.
-III-A2 Action Space
-Instead of directly predicting the next action of throttle, steer, and brake, we estimate the future waypoints $\mathcal{W}$ of the ego-vehicle in bird’s eye view space which is centered at the ego vehicle’s current coordinate frame.
-The trajectory consists of a sequence of 2D waypoints $\mathcal{W}=\{\bm{w_{t}}=(x_{t},y_{t})\}^{T}_{t=1}$.
-In our settings, $T=4$ which means the agent needs to predict 4 waypoints in the future for each frame.
-Then we use two fine-tuned PID controllers for lateral and longitudinal control to obtain steer, throttle, and brake values from the predicted waypoints.
-These two PID controllers are not involved in the training process.
-III-A3 Goal Location
-According to CARLA [22] 0.9.10’s protocol, the high-level goal locations $\bm{g}$ are provided as GPS coordinates.
-This goal location $\bm{g}$ is sparse (hundreds of meters apart) which can only be used as guidance.
-We kindly remind the readers that the goal location is different from those predicted waypoints, which are dense, and only a few meters away from each other.
-III-B Policy Learning
-In this paper, we plan to use the imitation learning mechanism to learn the goal-conditioned policy $\pi(\bm{a}|\bm{s},\bm{g})$.
-As we discussed previously, we intend to predict the future waypoints instead of directly learning the actions.
-An expert policy is applied in the environment to collect a dataset $\mathcal{D}=\{(\bm{s},\mathcal{W})\}^{N}$ with the size of $N$, which contains multiple tuples of the current observation $\bm{s}$ and a set of waypoints $\mathcal{W}$ in the future timesteps.
-Similar to previous works [7], [25], we also use $l_{1}$-norm based loss function because of its demonstrated robustness to outliers.
+3.2 Policy Learning
+In this paper, we use the imitation learning mechanism to learn the goal-conditioned policy $\pi(\bm{a}|\bm{s},\bm{g})$.
+As we discussed previously, the future waypoints instead of the actions are predicted.
+An expert policy is applied in the environment to collect a dataset $\mathcal{D}=\{(\bm{s},\mathcal{W})\}^{N}$ with the size of $N$, which contains tuples of the observation $\bm{s}$ and a set of waypoints $\mathcal{W}$ in future time steps.
+Similar to previous works [12], [7], we use $l_{1}$-norm based loss function because of its demonstrated robustness to outliers.
 For each input, the policy learning loss can be formalized as:
-$\mathcal{L}_{\textrm{pl}}=\sum_{t=1}^{T}||\bm{\hat{w}_{t}}-\bm{w_{t}}||_{1}$ (1)
+$\mathcal{L}_{\textrm{pl}}=\sum_{t=1}^{T}||\bm{\hat{w}_{t}}-\bm{w_{t}}||_{1}\text{,}$ (1)
 where $\bm{\hat{w}_{t}}$ is the $t$-th predicted waypoint and $\bm{w_{t}}$ is the t-th ground truth waypoint produced by the expert policy.
-$\mathcal{L}_{\textrm{pl}}$ is the loss function of policy learning.
-III-C Cross Semantics Generation
-The motivation for our approach is based on the fact that multi-modality inputs have common information and also their own unique information.
+[…]
+Figure 2: An Overview of Our Penalty-based Imitation Learning with Cross Semantics Generation.
+3.3 Contrastive-based Cross Semantics Generation
+Our approach’s motivation is based on the fact that multi-modality inputs have common information and also their own unique information.
 For instance, the vehicle’s and pedestrian’s shape and location are the shared information of LiDAR and camera input.
-The term unique information refers to information that is not present in other sources of input.
-Extracting and aligning the common information from both modalities, helps the decision network to find the logical connection between modalities and capture their global context since the extracted features are more reasonable and easier to understand.
-There are two problems we want to solve in this section.
-They are 1) How to find the shared information of these two modalities?
+Unique information refers to information that is not present in the other source of modality.
+Extracting and aligning the common information from both modalities can capture the global context within multi-modalities.
+In this section, we address two problems.
+1) How to find the shared information of these two modalities?
 2) How to align the extracted shared information in the same space?
 To address the first problem, we propose the cross semantics generation approach.
-We plan to utilize the LiDAR input to produce semantic information from the camera and also use the camera input to generate semantic information from the LiDAR.
-As Figure 1 demonstrates, the information flows for the LiDAR and camera semantic information generation are crossed since we used the information from one modality to generate the semantic segmentation of the other modality.
-That is the reason we name our approach cross semantics generation.
-The RGB image and the bird’s eye view of the LiDAR pseudo image are fed into two residual networks [26] to extract the features of these two modalities.
-We utilize two linear layers to extract the shared information of the camera and LiDAR.
-Those from the camera input extracted features are used to generate the top-down semantic segmentation which is aligned with the LiDAR input and those from the LiDAR input extracted features are used to generate the semantic segmentation of the front image.
-The reason we used the information to generate the semantic segmentation is that 1) semantic segmentation has less noise compared to the original data and 2) with the semantic segmentation we can filter out the unique information such as traffic light color and traffic sign pattern.
-In our setup, the semantic segmentation contains 4 channels, namely the drivable area, the non-drivable area, the object in the drivable areas (vehicle and pedestrians), and others.
+We utilize the LiDAR input to produce semantic segmentation of the camera input and use the camera input to generate semantic segmentation of the LiDAR pseudo image input.
+As Figure 2 demonstrates, the information flows of the LiDAR (gray) and camera (black) semantic information generation are crossed since we used the information from one modality to generate the semantic segmentation of the other modality.
+This is the reason that we name our approach cross semantics generation.
+The RGB image and the BEV of the LiDAR pseudo image are fed into two residual networks [18] to extract the features of these two modalities.
+We utilize two linear layers to extract the information shared by the camera and LiDAR.
+The feature embeddings are leveraged to generate the semantic segmentation because 1) semantic segmentation has less noise than the original data, and 2) with semantic segmentation, we filter out the unique information, such as traffic light color and traffic sign patterns.
+In our setup, the semantic segmentation contains 4 channels: the drivable area, the non-drivable area, the objects (vehicles and pedestrians) in the drivable areas, and others.
 We define $y_{\textrm{front}}$ as the ground truth front segmenation tensor with the shape of $160\times 768\times 4$ and $\hat{y}_{\textrm{front}}$ as the output of the front view decoder with the same shape.
@@
 The reconstruction losses can be written as $\mathcal{L}_{\textrm{front}}=\mathcal{L}_{\textrm{CE}}(\hat{y}_{\textrm{front}},y_{\textrm{front}})$ and $\mathcal{L}_{\textrm{td}}=\mathcal{L}_{\textrm{CE}}(\hat{y}_{\textrm{td}},y_{\textrm{td}})$, where $\mathcal{L}_{\textrm{CE}}(\cdot)$ indicates the cross entropy loss.
-$\mathcal{L}_{\textrm{front}}$ and $\mathcal{L}_{\textrm{td}}$ represent the loss function of front reconstruction loss and top-down reconstruction loss, respectively.
-To align the extracted shared information into the same space, we leverage the contrastive loss.
-The primary goal of contrastive loss is to learn an embedding space where similar examples are placed closer together, while dissimilar examples are placed farther apart.
+$\mathcal{L}_{\textrm{front}}$ and $\mathcal{L}_{\textrm{td}}$ represent the functions of front reconstruction loss and top-down reconstruction loss, respectively.
+To address the second problem, we leverage the contrastive loss to align the extracted shared information into the same space.
+The primary goal of contrastive loss is to learn an embedding space where similar examples are placed closer together while dissimilar examples are placed farther apart.
 This is achieved by defining a distance metric between the embeddings and then minimizing the distance between similar examples while maximizing the distance between dissimilar examples.
-Note that the embeddings extracted from LiDAR and camera inputs are mapped into two Gaussian distributions by generating mean and variance, using MLP networks.
-Suppose $P_{1}$ and $P_{2}$ represent two batches of Gaussian distributions, derived from RGB and LiDAR inputs.
+The embeddings extracted from LiDAR and camera inputs are mapped into two Gaussian distributions by generating mean and variance using MLP networks.
+Suppose $P_{1}$ and $P_{2}$ represent two batches of Gaussian distributions derived from RGB and LiDAR inputs.
 Each batch contains $N_{b}$ number of distributions.
 The contrastive loss can be formulated as follows:
-$\mathcal{L}_{\textrm{align}}({P_{1}},{P_{2}})=\bm{E}\odot\bm{D}+(1-\bm{E})\odot\max\{0,\epsilon_{\textrm{a}}-\bm{D}\}$ (2)
+$\mathcal{L}_{\textrm{align}}({P_{1}},{P_{2}})=mean(\bm{E}\odot\bm{D}+(1-\bm{E})\odot\max\{0,\epsilon_{\textrm{a}}-\bm{D}\})$ (2)
 where $\bm{D}$ is the matrix consisting of $d_{ij}=d(p_{1}^{i},p_{2}^{j})$.
@@
 $\odot$ represents the Hadamard product of two matrices.
-In our setting, we use the symmetric version of Kullback–Leibler divergence
-$d(p_{1}^{i},p_{2}^{j})=\frac{1}{2}\textrm{KL}(p_{1}^{i}||p_{2}^{j})+\frac{1}{2}\textrm{KL}(p_{2}^{j}||p_{1}^{i})$ (3)
-where $\textrm{KL}(\cdot)$ is the KL-divergence metric, as the metric to measure the difference between two distributions.
-The sampled shared embedding is concatenated with image and LiDAR embeddings, along with measurements (velocity, throttle, steer, and brake from the previous frame).
+$mean(\cdot)$ represents the mean value of one tensor.
+In our setting, we use the symmetric version of Kullback–Leibler divergence to measure the distance of two distributions:
+$d(p_{1}^{i},p_{2}^{j})=\frac{1}{2}\textrm{KL}(p_{1}^{i}||p_{2}^{j})+\frac{1}{2}\textrm{KL}(p_{2}^{j}||p_{1}^{i})\text{,}$ (3)
+where $\textrm{KL}(\cdot)$ is the KL-divergence metric as the metric to measure the difference between two distributions.
+The sampled shared embedding is concatenated with image and LiDAR embeddings, along with measurements (speed, throttle, steer, and brake from the previous frame).
 This combined information serves as the input for the penalty-based imitation learning module.
-III-D Auxiliary Tasks
-Auxiliary tasks have been proven to be efficient in many learning approaches.
-It has two main advantages.
-1) By introducing auxiliary tasks, guarantees the important information flows remain in the networks.
-This information flow is critical for the decision network to make decisions.
-2) Auxiliary tasks can guide the direction of gradient descent during training, which can lead the neural network to a more optimal location in the weight space.
-In this autonomous driving task, we would introduce two extra auxiliary tasks, namely traffic light classification and stop sign classification.
-III-D1 Traffic Light Classification
-The output of the traffic light decoder should be a vector of 4 which indicates these four states red light, yellow light, green light, and none in the current frame.
+3.4 Auxiliary Tasks
+Auxiliary tasks have been proven to be efficient in many learning approaches with two main advantages: 1) Guaranteeing the important information flows remain in the networks.
+This information flow is critical for the decision-making network.
+2) Guiding the direction of gradient descent during training, leading the neural network to a more optimal location in the weight space.
+In this autonomous driving task, we introduce two extra auxiliary tasks: traffic light classification and stop sign classification.
+3.4.1 Traffic Light Classification
+The output of the traffic light decoder should be a vector of 4, which indicates four states: red light, yellow light, green light, and none in the current frame.
 We then define $y_{\textrm{light}}$ as the ground truth traffic light vector of length 4 and $\hat{y}_{\textrm{light}}$ as the output of the traffic light decoder with the same shape.
 We use cross-entropy loss $\mathcal{L}_{\textrm{light}}=\mathcal{L}_{\textrm{CE}}(\hat{y}_{\textrm{light}},y_{\textrm{light}})$ for this task.
-III-D2 Stop Sign Classification
-The output of the stop sign decoder should have a vector of 2 which indicates if a stop sign exists in the current frame.
-The ground truth stop sign vector of length 2 and the output of the stop sign decoder with the same shape are defined as $y_{\textrm{stop}}$ and $\hat{y}_{\textrm{stop}}$, respectively.
-Also, cross-entropy loss $\mathcal{L}_{\textrm{stop}}=\mathcal{L}_{\textrm{CE}}(\hat{y}_{\textrm{stop}},y_{\textrm{stop}})$ is used for this task
+3.4.2 Stop Sign Classification
+The output of the stop sign decoder should have a vector of 1, indicating if a stop sign exists in the current frame.
+The ground truth stop sign vector of length 1 and the output of the stop sign decoder with the same shape are defined as $y_{\textrm{stop}}$ and $\hat{y}_{\textrm{stop}}$, respectively.
+Also, binary cross-entropy loss $\mathcal{L}_{\textrm{stop}}=\mathcal{L}_{\textrm{BCE}}(\hat{y}_{\textrm{stop}},y_{\textrm{stop}})$ is used for this task
 Note that these two tasks are trained simultaneously with policy generation networks and cross semantics generation tasks. In summary, the final loss function can be calculated as:
@@
 $\eta_{1}$, $\eta_{2}$, $\eta_{3}$, $\eta_{4}$, $\eta_{5}$ are weights to balance these losses.
-III-E Penalty-based Imitation Learning
-Our investigation revealed that the objective function used in imitation learning and the metric used for evaluating autonomous driving performance are not consistent with each other.
-Hence, achieving a low loss in the objective function does not necessarily guarantee a high driving score or successful route completion.
+3.5 Penalty-based Imitation Learning with Constraint Optimization
+Our investigation revealed that the objective function used in imitation learning and the metric used for evaluating autonomous driving performance are not consistent.
+Hence, achieving a low loss in the objective function does not necessarily guarantee a high driving score and route completion.
 We identified two possible factors contributing to this discrepancy.
   • The expert agent still makes mistakes when generating the dataset. Sometimes, the expert agent runs a red light and violates the stop sign rule.
-  • The objective function is not sensitive to serious violations of the traffic rules, i.e.
-the violation of red lights and stop signs.
-The average objective function loss may not increase too much when violating the traffic rules despite that this violation may cause serious consequences which result in a huge drop in driving score and route completion.
-Behavior Cloning (BC) is an imitation learning technique that seeks to replicate the behavior of an expert agent.
-As a result, the performance of the trained agent is inherently limited by the capabilities of the expert agent.
-In other words, if the expert agent makes a mistake, the trained agent will learn to replicate that mistake rather than correct it.
-Our goal is to modify the objective function used in imitation learning to incorporate traffic rules.
-Specifically, we plan to introduce additional penalties (higher loss), when the agent generates short-term future waypoints that violate traffic rules during training.
-Traffic rules can be expressed as constraint functions that define conditions the optimization problem must meet.
-In our approach, we focus on three specific aspects of driving behavior: running red lights, ignoring stop signs, and failing to slow down when turning.
+  • The objective function is not sensitive to serious violations of the traffic rules, i.e., the violation of red lights and stop signs.
+The average objective function loss may not increase too much when violating the traffic rules, although this violation may cause serious consequences, which result in a huge drop in driving score and route completion.
+We aim to constrain the objective function of imitation learning to incorporate traffic rules.
+Traffic rules can be represented as constraint functions that specify the conditions the optimization problem must satisfy.
+Our approach focuses on three specific aspects: running red lights, ignoring stop signs, and failing to slow down when turning.
 These are the primary issues we observed with our vanilla imitation learning method.
-To address these concerns, we propose three corresponding penalties that can be used to quantify and penalize these violations.
-III-E1 Red Light Penalty
+To address these issues, we propose three corresponding penalties that can be used to quantify and penalize these violations.
+3.5.1 Red Light Penalty
 For the red light violation, we design a red light penalty as follows:
-${\mathcal{P}}_{\rm tl}=\mathbb{E}_{(\bm{s},\mathcal{W})\sim\mathcal{D}}[\mathbb{1}_{\rm red}\cdot\sum_{t=1}^{T}c_{i}\cdot{\rm max}\{0,y_{t}-\overline{d}\}]\text{,}$ (5)
-where $y_{t}$ represents the $y$ axis value in the $t$-th predicted waypoints $\bm{\hat{w}_{t}}$ generated by the trained agent and $\overline{d}$ denotes the distance between the ego car and the stop line at the intersection along the forward direction of the ego car ($y$ axis).
+${\mathcal{P}}_{\rm tl}=\mathbb{1}_{\rm red}\cdot\sum_{t=1}^{T}c_{i}\cdot{\rm max}\{0,\hat{y}_{t}-y_{\textrm{stop}}\}\text{,}$ (5)
+where $\hat{y}_{t}$ represents the $y$ axis value in the $t$-th predicted waypoints $\bm{\hat{w}_{t}}=(\hat{x}_{t},\hat{y}_{t})$ and $y_{\textrm{stop}}$ denotes the distance between the ego car and the stop line at the intersection along the forward direction of the ego car ($y$ axis).
 The weight parameter is denoted by $c_{i}$, and the sum of all weight parameters is equal to one.
 $\mathbb{1}_{\rm red}$ indicates the presence of a red light that may affect the agent in the current frame.
-$(\bm{s},\mathcal{W})$ represents the combination of observation state $\bm{s}$ and waypoints set $\mathcal{W}$ of one frame, while $\mathcal{D}$ refers to the entire dataset.
 When facing red lights, a red light penalty is added based on the distances of the predicted waypoints beyond the stop line at the intersection.
@@
 The necessary information for calculating the red light penalty, such as traffic light information and stop line location, is pre-processed and stored in each frame of our dataset.
-III-E2 Stop Sign Penalty
+3.5.2 Stop Sign Penalty
 Similar to the red light penalty, a stop sign penalty is given when the predicted waypoints violate the stop sign rule. The penalty is formalized as follows:
-${\mathcal{P}}_{\rm ss}=\mathbb{E}_{(\bm{s},\mathcal{W})\sim\mathcal{D}}[\mathbb{1}_{\rm stopsign}\cdot{\rm max}\{v-\epsilon_{v},0\}]$ (6)
+${\mathcal{P}}_{\rm ss}=\mathbb{1}_{\rm stopsign}\cdot{\rm max}\{v-\epsilon_{v},0\}$ (6)
 where $v$ is the estimated speed calculated by
 $v=\frac{||\bm{\hat{w}_{0}}-\bm{\hat{w}_{1}}||_{2}}{\Delta t}$ (7)
-The variables $\bm{\hat{w}_{0}}$ and $\bm{\hat{w}_{1}}$ represent the first and second predicted waypoints, respectively, while $\Delta t$ indicates the time interval between each frame during data collection.
+The variables $\bm{\hat{w}_{0}}$ and $\bm{\hat{w}_{1}}$ represent the first and second predicted waypoints, respectively, while $\Delta t$ indicates the time interval between each frame.
 The function $\mathbb{1}_{\rm stopsign}$ serves as an indicator for stop sign checking.
-The maximum speed required to pass stop sign tests is denoted by $\epsilon_{v}$.
-An upper-speed limit is established for an area affected by a stop sign, and only speeds lower than this limit are permitted for the agent to pass through.
+The maximum speed to pass stop sign tests is denoted by $\epsilon_{v}$.
+An upper-speed limit $\epsilon_{v}$ (close to zero) is established for an area affected by a stop sign, and only speeds lower than this limit are permitted for the agent to pass through.
 If the agent exceeds this speed limit, a penalty is imposed based on its speed.
-As the training network only generates predicted waypoints, the velocity estimated from the waypoints is used to compute the stop sign penalty.
-III-E3 Speed Penalty
-If the agent attempts to turn at excessive speed, a penalty will be enforced.
-The rationale behind this penalty is based on human driving experience, as it is commonly known that turning at high speeds can increase the risk of collisions with pedestrians or other objects due to reduced braking distance.
+As the training network only generates predicted waypoints, the speed estimated from the waypoints is used to compute the stop sign penalty.
+3.5.3 Curvature Speed Penalty
+A penalty will be enforced if the agent attempts to turn at excessive speed.
+The rationale behind this penalty is based on human driving experience, as it is commonly known that turning at high speeds can increase the risk of collisions with pedestrians or other objects due to longer braking distance.
 The speed penalty is defined as follows:
-${\mathcal{P}}_{\rm sp}=\mathbb{E}_{(\bm{s},\mathcal{W})\sim\mathcal{D}}[{\rm\sin}(\Delta\delta)\cdot{\max}\{v-v_{\textrm{lb}},0\}]$ (8)
+${\mathcal{P}}_{\rm sp}={\rm\sin}(\Delta\delta)\cdot{\max}\{v-v_{\textrm{lb}},0\}$ (8)
 where $\Delta\delta$ denotes the deviation in direction between the current frame and the next frame.
@@
 Any speed below this limit is not subject to the speed penalty.
-III-E4 Objctive Function
-We trained the network to minimize the difference between the predicted waypoints and the ground truth waypoints, applying an $l_{1}$-norm based loss function.
+3.5.4 Objctive Function
+We trained the network to minimize the difference between the predicted and ground truth waypoints, applying an $l_{1}$-norm-based loss function.
 Also, cross semantics generation and auxiliary tasks are trained together with the policy generation task.
-The final loss function is defined in Equation (4).
+Equation (4) defines the final loss function.
 By applying the penalties, we formalize the constrained optimization:
 By applying the penalties, we formalize the constrained optimization:
-$\min_{\bm{\Theta}}\quad$ $\mathcal{L}_{\textrm{final}}$ (9)
+$\min_{\bm{\Theta}}$ $\mathcal{L}_{\textrm{final}}$ (9)
 s.t. $\mathcal{P}_{\rm tl},\mathcal{P}_{\rm ss},\mathcal{P}_{\rm sp}=0$
@@
 Well-chosen $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$ are important for optimization.
-According to our experiments, too large hyper-parameters influence the behaviors in other scenarios while too small ones are not powerful enough for the agent to obey the corresponding traffic rules.
+According to our experiments, hyperparameters that are too large influence behaviors in other scenarios while hyperparameters that are too small are not powerful enough for the agent to obey the corresponding traffic rules.
 The right part of Figure 2 demonstrates the process of penalty-based imitation learning in detail.
-IV Experiments
-IV-A Task Description
-We focus on a navigation task where the vehicle navigates along predefined routes in various scenarios and areas.
-The vehicle is guided by GPS signals and situations where GPS signals are low or not available are not being considered.
-The predefined routes include specific scenarios to test the agent’s ability to respond to emergencies such as obstacle avoidance, other vehicles running red lights, and the sudden appearance of pedestrians on the road.
-The agent is required to complete the route within a specified time limit, and any time exceeding the limit is considered a failure in terms of route completion.
-IV-B Training Dataset
-Obtaining realistic driving data is a challenging task.
-Therefore, we opted to use the Carla simulator [22] to gather training data that has been processed by the expert policy.
-Our training dataset is the same as TransFuser [6], which comprises approximately 2500 routes through junctions in 8 different towns.
-These routes have an average length of 100 meters, and around 1000 routes along curved highways with an average length of 400 meters.
-We utilized the expert agent, similar to TransFuser, to generate this training data.
-The expert policy we use is the same export policy in TransFuser [6].
+4 Experiments
+4.1 Training Dataset
+Obtaining realistic driving data is challenging.
+Therefore, we opted to use the CARLA simulator to gather training data that has been processed by the expert policy.
+Our training dataset comprises approximately 2500 routes through junctions in 8 different towns.
+These routes have an average length of 100 meters and around 1000 routes along curved highways with an average length of 400 meters.
+The expert policy we use is the updated one in TransFuser+.
 For each frame, We collect
@@
   • The cloud point of LiDAR with 180 degrees in front of the vehicle and the top-down semantic segmentation.
-  • The measurements of the steer, throttle, brake action values, and velocity in the current frame.
-  • The traffic light information which includes the stop line position, and traffic light status.
+  • The steer, throttle, brake action values, and speed measurements in the current frame.
+  • The traffic light information, which includes the stop line position and traffic light status.
   • The stop sign information indicates if the current frame is influenced by a stop sign.
   • The future waypoints in the next 4 frames.
-Note that the interval for each frame is 0.5 seconds, indicating 2FPS collection frequency.
-IV-C Metrics
-The metrics used to evaluate the behavior of each are Route completion and Infraction Scores.
-IV-C1 Route Completion
-Route completion (RC) refers to the proportion of the completed route out of the whole route.
+Note that the interval for each frame is 0.5 seconds, indicating 2FPS frequency.
+4.2 Metrics
+Route completion and infraction scores are used to evaluate the agent’s behavior.
+Route Completion Route completion (RC) refers to the proportion of the completed route out of the whole route.
 Suppose $R_{j}$ means the route completion proportion in route $j$ and $N_{\textrm{route}}$ means the number of routes.
@@
 $RC=\frac{1}{N_{\textrm{route}}}\sum_{j}^{N_{\textrm{route}}}R_{j}$ (11)
-IV-C2 Infraction Score
-Infraction Score (IS) is used to measure the driving behavior of the agent.
-We define $p_{k}$ as the penalty for an infraction instance, $k$ is incurred by the agent and $n_{k}$ is the number of occurrences of infraction instance $k$.
-Then the Infraction Score can be defined by:
-$IS=\prod_{k}^{\{\rm Ped,Veh,Stat,Red,Stop\}}s_{k}^{n_{k}}$ (12)
+Infraction Score Infraction Score (IS) is used to measure the driving behavior of the agent.
+We define $p_{k}$ as the penalty for an infraction instance, $k$ is incurred by the agent, and $n_{k}$ is the number of occurrences of infraction instance $k$.
+Then, the infraction score can be defined by:
+$IS=\prod_{k}^{\{\rm Ped,Veh,Stat,Red,Stop\}}s_{k}^{n_{k}}\text{.}$ (12)
 The infraction instances include collision with a pedestrian, collision with a vehicle, collision with static layout, red light violations, and stop sign violations.
@@
 $n_{k}$ is the number of infractions.
-IV-C3 Driving Score
-Driving Score (DS) aims to measure the overall driving performance. It is defined by the weighted average of the route completion with an infraction multiplier:
-$DS=\frac{1}{N_{\textrm{route}}}\sum_{j}^{N_{\textrm{route}}}R_{j}IS_{j}$ (13)
-where $R_{j}$ and $IS_{j}$ are the route completion and infraction score $IS$ (12) for $j$-th test route.
-IV-C4 Infractions per km
-Infractions per km refer to the average infraction numbers per kilometer.
-The infractions include collisions with pedestrians, vehicles, and static elements, running a red light, off-road infractions, route deviations, timeouts, and vehicle blocks.
-${\rm Infractions\ per\ km}=\frac{\sum_{j}^{N_{\textrm{route}}}\#\ {\rm infractions}_{j}}{\sum_{j}^{N_{\textrm{route}}}l_{j}}$ (14)
-where $l_{j}$ is the driven distance (in kilometers) for route $j$.
-IV-D Test Results
-We use Carla leaderboard - Town05 long benchmarks to evaluate our model.
-Town05 long benchmark contains 10 routes and all of these routes are over 2.5 km.
-This benchmark is also used by InterFuser [8] and TransFuser [21].
-We also test our model in the Longest6 benchmark which is proposed in TransFuser+ [7].
-Baseline.
-The other baselines we chose to compare with our model are TransFuser+, TransFuser, Geometric Fusion, and LateFusion.
-TransFuser [21] introduces the Transformer into the multi-sensor fusion architecture to achieve better end-to-end autonomous driving results.
-TransFuser+ [7], as an extension of TransFuser, leverages several auxiliary losses to ensure important information flows such as traffic light and road line information in the network.
-InterFuser [8] developed a safety control module to regulate the behaviors of the agent, preventing the agent from violating the traffic rules.
-LAV [9] utilizes data from both the ego vehicle and surrounding vehicles to enhance the dataset through data augmentation.
-This is achieved by learning a spatial intermediate representation that remains invariant to different viewpoints.
-The evaluation results demonstrated in Table I indicate that our model has a huge increase in driving scores and infraction penalties compared to other SOTA models.
-As the table demonstrates, our model adheres to traffic rules more effectively in terms of collision, red light violations, and stop sign violations.
-Table II demonstrates the performance of our model in the longest6 benchmark provided by Transfuser+.
-Our model continues to surpass all other baselines in driving scores and infraction scores.
-Additionally, there has been a significant decrease in collisions with pedestrians and vehicles.
-While our model may not attain the top route completion score, its highest infraction score and driving score suggest that the model embraces a more cautious approach to mitigating traffic rule violations, ultimately leading to the highest driving score.
-IV-E Ablation Studies
+Driving Score Driving Score (DS) aims to measure the overall driving performance. It is defined by the weighted average of the route completion with an infraction multiplier:
+$DS=\frac{1}{N_{\textrm{route}}}\sum_{j}^{N_{\textrm{route}}}R_{j}IS_{j}\text{,}$ (13)
+where $R_{j}$ and $IS_{j}$ are route completion, infraction score $IS$ (12) for $j$-th route.
+4.3 Test Results
+We use CARLA Leaderboard - Town05 Long Benchmark and Longest6 Benchmark to evaluate our model.
+Town05 Long Benchmark contains 10 routes, all of which are over 2.5 km.
+This benchmark is also used by InterFuser, and TransFuser.
+Longest6 benchmark is proposed in TransFuser+ with increased traffic density, and challenging pre-crash traffic scenarios.
+Model Driving score Route compl. Infrac. score
+[…]
+P-CSG(Ours) $\textbf{59.10}\pm 4.07$ $86.84\pm 2.99$ $\textbf{0.68}\pm 0.06$
+TransFuser [29] $34.50\pm 2.54$ $61.16\pm 4.75$ $0.56\pm 0.06$
+TransFuser+ [12] $36.19\pm 0.90$ $70.13\pm 6.80$ $0.51\pm 0.03$
+InterFuser [33] $50.64\pm 3.51$ $89.13\pm 4.12$ $0.57\pm 0.05$
+LAV [6] $45.20\pm 6.36$ $\textbf{91.55}\pm 5.61$ $0.49\pm 0.06$
+Table 1: CARLA Leaderboard - Town 05 Long Benchmark.
+Model Driving score Route compl. Infrac. score
+[…]
+P-CSG(Ours) $\textbf{49.21}\pm 2.93$ $86.24\pm 2.46$ $\textbf{0.56}\pm 0.04$
+WOR [5] $20.53\pm 3.12$ $48.47\pm 3.86$ $0.56\pm 0.03$
+LAV [6] $32.74\pm 1.45$ $70.36\pm 3.14$ $0.51\pm 0.02$
+Late Fusion [35] $22.47\pm 3.71$ $83.30\pm 3.04$ $0.27\pm 0.04$
+TransFuser+ [12] $47.30\pm 5.72$ $\textbf{93.38}\pm 1.20$ $0.50\pm 0.06$
+Table 2: CARLA Leaderboard - Longest6 Benchmark.
+The evaluation results in Table 2 indicate that our model significantly increases driving scores and infraction penalties compared to other baselines.
+Table 2 demonstrates the performance of our model in the Longest6 Benchmark provided by Transfuser+.
+Our model surpasses all other baselines in driving and infraction scores.
+While our model may not attain the top route completion score, its highest infraction score and driving score suggest that it embraces a more cautious approach to mitigating traffic rule violations, ultimately leading to the highest driving score.
+4.4 Ablation Studies
 In this subsection, we analyze the influences of different penalty weights for corresponding traffic rules.
-As Table III demonstrates, two extra weights for each penalty are selected for comparison.
+The ablation study focuses on significant hyper-parameters, namely $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$, which correspond to the penalty weights associated with the red light, speed, and stop sign.
+Our optimal model employs default weights of $\lambda_{1}=0.5$, $\lambda_{2}=0.05$, and $\lambda_{3}=0.5$.
+As Table 3 demonstrates,
+Model Driving score Route compl. Infrac. score
+[…]
+P-CSG(Ours) $\textbf{59.10}\pm 4.07$ $86.84\pm 2.99$ $\textbf{0.68}\pm 0.06$
+No CSG $45.67\pm 5.51$ $84.82\pm 2.25$ $0.52\pm 0.10$
+No Penalty $37.32\pm 3.58$ $78.41\pm 11.32$ $0.54\pm 0.21$
+$\lambda_{1}=0.7$ $54.36\pm 1.78$ $84.37\pm 5.01$ $0.64\pm 0.04$
+$\lambda_{1}=0.3$ $56.80\pm 2.22$ $89.60\pm 1.90$ $0.60\pm 0.02$
+$\lambda_{2}=0.07$ $51.54\pm 1.50$ $81.49\pm 1.42$ $0.63\pm 0.02$
+$\lambda_{2}=0.03$ $54.59\pm 2.88$ $88.12\pm 1.95$ $0.64\pm 0.02$
+$\lambda_{3}=0.7$ $55.98\pm 3.00$ $\textbf{89.81}\pm 3.33$ $0.63\pm 0.02$
+$\lambda_{3}=0.3$ $51.98\pm 1.93$ $84.66\pm 3.53$ $0.63\pm 0.01$
+Table 3: Ablation Study.
+two extra weights for each penalty are selected for comparison.
 We also provide the results of models without CSG and penalties for comprehensive analysis.
-We notice that the infractions of traffic lights and stop signs are largely reduced by adding penalties.
-Our proposed multi-sensor fusion technology (CSG) also decreases the possibility of hitting obstacles such as vehicles, pedestrians, and other statics.
-The results of different penalty weights are also listed in the table for comparison.
+The results of different penalty weights are also listed in the table.
 We found that assigning greater weight to more severe violations will increase the performance of our model.
-For instance, we apply greater penalties for the red light and the stop sign violations compared to overspeeding by turning, since those two violations cause more serious consequences.
-Model Driving score Route compl. Infrac. score Collision pedestrian Collision vehicle Collision static Red light infraction Stop Sign Infraction
-[…]
-P-CSG(Ours) $\textbf{59.10}\pm 4.07$ $86.84\pm 2.99$ $\textbf{0.68}\pm 0.06$ $0.01\pm 0.01$ $\textbf{0.07}\pm 0.05$ $0\textbf{0.00}\pm 0$ 0.00 $0.04\pm 0.02$ $0\textbf{0.00}\pm 0$ 0.00
-TransFuser [21] $34.50\pm 2.54$ $61.16\pm 4.75$ $0.56\pm 0.06$ $0.01\pm 0.01$ $0.58\pm 0.07$ $0.38\pm 0.05$ $0.12\pm 0.03$ $0.05\pm 0.02$
-TransFuser+ [7] $36.19\pm 0.90$ $70.13\pm 6.80$ $0.51\pm 0.03$ $0\textbf{0.00}\pm 0$ 0.00 $0.40\pm 0.13$ $0.04\pm 0.03$ $0.11\pm 0.10$ $0.04\pm 0.01$
-InterFuser [8] $50.64\pm 3.51$ $89.13\pm 4.12$ $0.57\pm 0.05$ $0\textbf{0.00}\pm 0$ 0.00 $0.09\pm 0.04$ $0.01\pm 0.01$ $\textbf{0.02}\pm 0.01$ $0.04\pm 0.01$
-LAV [9] $45.20\pm 6.36$ $\textbf{91.55}\pm 5.61$ $0.49\pm 0.06$ $0\textbf{0.00}\pm 0$ 0.00 $0.92\pm 0.42$ $0.33\pm 0.50$ $0.28\pm 0.28$
-TABLE I: Carla Leaderboard - Town 05 Long Benchmark.
-In our assessment of the Town05 long benchmark, we analyze several performance indicators including the Driving Score, Route Complication, Infraction Score, instances of collisions, violations of red light signals, and violations related to stop signs.
-These metrics are precisely outlined by the Carla leaderboard [22].
-It’s important to highlight that all alternate baseline models undergo retraining and testing within our dedicated CARLA environment.
-Each model’s evaluation is conducted thrice, employing distinct random seeds on every occasion.
-The performance data of LAV is collected from their paper.
-Model Driving score Route compl. Infrac. score Collision pedestrian Collision vehicle Collision static Red light infraction Stop Sign Infraction
-[…]
-P-CSG(Ours) $\textbf{49.21}\pm 2.93$ $86.24\pm 2.46$ $\textbf{0.56}\pm 0.04$ 0.00 0.18 0.24 0.21 0.00
-WOR [27] $20.53\pm 3.12$ $48.47\pm 3.86$ $0.56\pm 0.03$ 0.18 1.05 0.37 1.28 -
-LAV [9] $32.74\pm 1.45$ $70.36\pm 3.14$ $0.51\pm 0.02$ 0.16 0.83 0.15 0.96 -
-Late Fusion $22.47\pm 3.71$ $83.30\pm 3.04$ $0.27\pm 0.04$ 0.05 4.63 0.28 0.11 -
-TransFuser+ [7] $47.30\pm 5.72$ $\textbf{93.38}\pm 1.20$ $0.50\pm 0.06$ 0.03 2.45 0.07 0.16 -
-TABLE II: Carla Leaderboard - Longest6 Benchmark We assess our model with Longest6 benchmark provided by Transfuser+ [7].
-This benchmark is also under the Carla leaderboard framework.
-The metrics of the driving score, route complication, infraction score, collisions, violations of red light signals, and violations related to stop signs are the same as the metrics in the Carla leaderboard.
-The performance data of all other baselines is collected in their paper.
-Model Driving score Route compl. Infrac. score Collision pedestrian Collision vehicle Collision static Red light infraction Stop Sign Infraction
-[…]
-P-CSG(Ours) $\textbf{59.10}\pm 4.07$ $86.84\pm 2.99$ $\textbf{0.68}\pm 0.06$ $0.01\pm 0.01$ $\textbf{0.07}\pm 0.05$ $0\textbf{0.00}\pm 0$ 0.00 $0.09\pm 0.02$ $0\textbf{0.00}\pm 0$ 0.00
-No CSG $45.67\pm 5.51$ $84.82\pm 2.25$ $0.52\pm 0.10$ $0.02\pm 0.01$ $0.14\pm 0.03$ $0.03\pm 0.02$ $0.09\pm 0.02$ $0.01\pm 0.01$
-No Penalty $37.32\pm 3.58$ $78.41\pm 11.32$ $0.54\pm 0.21$ $0.02\pm 0.01$ $0.11\pm 0.03$ $0.02\pm 0.01$ $0.11\pm 0.03$ $0.04\pm 0.02$
-$\lambda_{1}=0.7$ $54.36\pm 1.78$ $84.37\pm 5.01$ $0.64\pm 0.04$ $0.01\pm 0.01$ $0.18\pm 0.02$ $0\textbf{0.00}\pm 0$ 0.00 $\textbf{0.08}\pm 0.01$ $0.01\pm 0.00$
-$\lambda_{1}=0.3$ $56.80\pm 2.22$ $89.60\pm 1.90$ $0.60\pm 0.02$ $0.01\pm 0.01$ $0.18\pm 0.01$ $0\textbf{0.00}\pm 0$ 0.00 $0.13\pm 0.01$ $0.01\pm 0.01$
-$\lambda_{2}=0.07$ $51.54\pm 1.50$ $81.49\pm 1.42$ $0.63\pm 0.02$ $0.02\pm 0.01$ $0.13\pm 0.10$ $0\textbf{0.00}\pm 0$ 0.00 $0.13\pm 0.01$ $0.01\pm 0.01$
-$\lambda_{2}=0.03$ $54.59\pm 2.88$ $88.12\pm 1.95$ $0.64\pm 0.02$ $0.01\pm 0.01$ $0.11\pm 0.01$ $0\textbf{0.00}\pm 0$ 0.00 $\textbf{0.08}\pm 0.06$ $0.01\pm 0.01$
-$\lambda_{3}=0.7$ $55.98\pm 3.00$ $\textbf{89.81}\pm 3.33$ $0.63\pm 0.02$ $0.01\pm 0.01$ $0.09\pm 0.02$ $0\textbf{0.00}\pm 0$ 0.00 $0.09\pm 0.04$ $0\textbf{0.00}\pm 0$ 0.00
-$\lambda_{3}=0.3$ $51.98\pm 1.93$ $84.66\pm 3.53$ $0.63\pm 0.01$ $0\textbf{0.00}\pm 0$ 0.00 $0.08\pm 0.03$ $0.02\pm 0.00$ $0.10\pm 0.02$ $0.01\pm 0.01$
-TABLE III: Ablation Study - Town 05 Long Benchmark The ablation study focuses on significant hyper-parameters, namely $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$, which correspond to the penalty weights associated with the red light, speed, and stop sign, respectively.
-Each model is evaluated three times with random seeds.
-Our optimal model employs default weights of $\lambda_{1}=0.5$, $\lambda_{2}=0.05$, and $\lambda_{3}=0.5$.
-During the ablation study, we isolate the impact of a single hyper-parameter by modifying it while keeping the others constant.
-V Robustness Study
-Attack Model Driving score Route compl. Infrac. score Collision pedestrian Collision vehicle Collision static Red light infraction Stop Sign Infraction
-[…]
-FGSM Attack $(\epsilon=0.01)$ P-CSG(Ours) $\mathbf{12.80}\pm 1.35$ $21.70\pm 2.87$ $\mathbf{0.59}\pm 0.02$ $\mathbf{0.00}\pm 0.00$ $\mathbf{0.22}\pm 0.23$ $2.40\pm 1.00$ $\mathbf{0.07}\pm 0.11$ $\mathbf{0.00}\pm 0.00$
-Transfuser+ $2.69\pm 0.73$ $42.44\pm 3.65$ $0.22\pm 0.06$ $\mathbf{0.00}\pm 0.00$ $1.23\pm 0.31$ $0.26\pm 0.17$ $0.08\pm 0.03$ $0.04\pm 0.02$
-Interfuser* $8.37\pm 1.06$ $\mathbf{72.04}\pm 4.15$ $0.15\pm 0.04$ $0.00\pm 0.01$ $0.64\pm 0.27$ $\mathbf{0.10}\pm 0.03$ $0.22\pm 0.03$ $0.07\pm 0.01$
-Dot Attack P-CSG(Ours) $\mathbf{36.60}\pm 3.17$ $\mathbf{75.53}\pm 6.43$ $\mathbf{0.56}\pm 0.07$ $0.01\pm 0.02$ $\mathbf{0.06}\pm 0.03$ $\mathbf{0.00}\pm 0.00$ $\mathbf{0.10}\pm 0.04$ $\mathbf{0.01}\pm 0.00$
-Transfuser+ $0.53\pm 0.21$ $54.06\pm 5.48$ $0.02\pm 0.01$ $\mathbf{0.00}\pm 0.00$ $2.53\pm 0.72$ $0.99\pm 0.16$ $0.21\pm 0.04$ $0.04\pm 0.01$
-Interfuser* $12.56\pm 2.15$ $73.62\pm 7.32$ $0.19\pm 0.02$ $0.01\pm 0.01$ $0.68\pm 0.26$ $\mathbf{0.00}\pm 0.00$ $0.21\pm 0.02$ $0.09\pm 0.02$
-TABLE IV: Attack Results.
-This table presents a comparative analysis of the performance and robustness of autonomous driving models under two types of adversarial attacks: FGSM [28] and Dot Attacks [29].
-Driving Score serves as the primary metric of overall performance, while the infraction score aggregates various safety-related factors.
-P-CSG, our proposed model, consistently outperforms the other models across these metrics.
-Interfuser* has the same PID Controller as P-CSG and Transfuser+ to ensure a fair comparison.
-Given the life-critical nature of autonomous driving, it becomes essential to examine the robustness of neural network models against such vulnerabilities.
-In this regard, we conduct a comprehensive comparative analysis to evaluate the robustness of our model.
-We benchmark its performance against two state-of-the-art architectures—Transfuser+ [7] and Interfuser [8].
-We focus on two specific types of white-box sensor attacks.
-These are Fast Gradient Sign Method (FGSM) Attack [28] and Dot Attack [29].
-FGSM Attack [28]: FGSM Attack is notable for their computational efficiency and minimal perceptibility to the human eye.
-This approach exploits the gradient of the model’s loss function with respect to the input data.
-$\bm{\eta}=\epsilon_{\textrm{p}}\operatorname{sign}(\nabla_{I}J(\bm{\Theta},\bm{I},y)).$ (15)
-$\bm{\tilde{I}}=\bm{I}+\bm{\eta}$ (16)
-where $\bm{I}$ is the original image input, and $y$ is the ground true label.
-The loss function is $J(\bm{\Theta},\bm{I},y)$, and its gradient with respect to ${\bm{I}}$ is $\nabla_{\bm{I}}J(\bm{\Theta},\bm{I},y)$.
-The hyperparameter $\epsilon_{\textrm{p}}$ controls perturbation magnitude, leading to adversarial input $\bm{\tilde{{I}}}$.
-$\operatorname{sign}(\cdot)$ is the sign indicator which has the value of 1 for positive value and -1 for negative value.
-The method creates subtle but impactful perturbations in autonomous driving.
+For instance, we apply greater penalties for the red light and the stop sign violations than overspeeding by turning since those two violations cause more serious consequences.
+5 Robustness Study
+Refer to caption(a)Refer to caption(b)Refer to caption(c)
 Figure 3: Qualitative Attack Results on P-CSG.
@@
 The subtle FGSM perturbation in (c) is hard to spot compared to (a), showcasing the method’s ability to create changes that are invisible to the human eye but still mislead models.
-Dot Attack[29] : Dot Attack targets deep learning algorithms by introducing minor, physical disruptions to camera systems.
-Specifically, a sticker with a designed dot pattern is placed on the camera lens.
-Given an RGB image input $\bm{I}$, semi-transparent dots are strategically placed.
-We first consider the perturbation function for a single dot on the image $\bm{I}$.
-The perturbed image $\bm{\tilde{I}}$ at pixel location $(x_{i},x_{j})$ can be written as
-$\begin{split}\bm{\tilde{I}}[x_{i},x_{j}]&=f_{0}(\bm{\phi_{0}};x_{i},x_{j})\\ &=(1-h(x_{i},x_{j}))\cdot\bm{I}(x_{i},x_{j})+h(x_{i},x_{j})\cdot\gamma\text{,}\end{split}$ (17)
-where
-$h(x_{i},x_{j})$ $=\alpha\cdot\exp\left(-\left(\frac{(x_{i}-x_{i}^{(c)})^{2}+(x_{j}-x_{j}^{(c)})^{2}}{r^{2}}\right)^{\beta}\right)$ (18)
-$\bm{\phi_{0}}$ $=(\gamma,(x_{i}^{(c)},x_{j}^{(c)}),r,\alpha,\beta)\text{.}$ (19)
-Here, $h(x_{i},x_{j})$ quantifies the balance between the original color at pixel $\bm{I}[x_{i},x_{j}]$ and a learnable color $\gamma$.
-This balance reduces with increasing distance from the center of the dot $\left(x_{i}^{(c)},x_{j}^{(c)}\right)$.
-$\bm{\phi_{0}}$ denotes the learnable parameters.
-By composing $N_{\textrm{dot}}$ of these single-dot perturbations:
-$f(\bm{\phi_{0}},\bm{\phi_{1}},...,\bm{\phi_{N_{\textrm{dot}}}};x_{i},x_{j})=f(\bm{\phi_{0}};x_{i},x_{j})\circ...\circ f(\bm{\phi_{N_{\textrm{dot}}}};x_{i},x_{j})\text{,}$ (20)
-the pixel at the location $(x_{i},x_{j})$ of the final perturbed image $\bm{\tilde{I}}$ can be formalized as:
-$\bm{\tilde{I}}[x_{i},x_{j}]=f_{0}(\bm{\phi_{0}},\bm{\phi_{1}},...,\bm{\phi_{N_{\textrm{dot}}}};x_{i},x_{j})\text{.}$ (21)
-These parameters are learned by increasing the loss of the model, leading to poor performance of the trained model.
-In Fig.
-2(b), the resulting dots mimic real-world lens stains, displaying a color gradient.
+Given the life-critical nature of autonomous driving, examining the robustness of neural network models against such vulnerabilities becomes essential.
+In this regard, we conduct a comprehensive comparative analysis to evaluate the robustness of our model.
+We benchmark its performance against two state-of-the-art methods — Transfuser+ and Interfuser.
+We focus on two specific types of white-box sensor attacks.
+These are Fast Gradient Sign Method (FGSM) Attack and Dot Attack.
+[^1] Interfuser* has the same PID Controller as P-CSG and Transfuser+ to ensure a fair comparison.
+  • FGSM Attack is notable for its computational efficiency and minimal perceptibility to the human eye.
+It uses the gradient of the model’s loss function to slightly alter the input image, creating an adversarial image that looks almost identical but can mislead the model.
+  • Dot Attack uses special stickers with dot patterns on camera lenses to subtly blur parts of the image.
+These near-invisible dots mislead deep learning models without being noticeable to humans.
+Each dot’s color blends the original and a new RGB value, decreasing transparency from the center outward.
+This technique effectively disrupts model performance with minimal visual impact.
 Unlike computational attacks like FGSM, Dot Attacks are easily deployed in the real world, requiring only the application of a sticker to the camera lens.
-V-A Experimental Design
-In our experiments conducted within the CARLA simulation environment, we utilized a predefined trajectory, termed Town05 Long.
-To ensure a fair comparison across three distinct deep learning models—Transfuser+, Interfuser, and P-CSG—a consistent PID Controller was applied to all models.
-This uniform approach was designed to specifically assess the inherent robustness of each model under different adversarial attacks, by eliminating variations in control mechanisms.
-FGSM Attacks: Our goal is to prompt immediate deviations in the vehicle’s trajectory by assigning target waypoints for the next four timesteps at $(-1.0,-2.0)$, $(-2.0,-4.0)$, $(-3.0,-6.0)$, and $(-4.0,-8.0)$, relative to the vehicle’s current position.
-These are used to formulate a dedicated loss function, from which adversarial perturbations are generated and applied to the input RGB images.
-We set the perturbation magnitude $\epsilon$ to 0.01 for these experiments.
-Dot Attacks: Unlike FGSM Attacks, which necessitate per-frame computations, Dot Attacks employ a pre-trained, universally applicable Dot Mask.
-Our experimental procedure initiated with the nine-dots mask training phase, followed by its application within the CARLA simulation framework.
-This structured approach permitted us to evaluate the practicality and effectiveness of Dot Attack under conditions that closely mimic real-world scenarios.
-V-B Results
+5.1 Results
 In our evaluation, the infraction score serves as the primary metric, encapsulating various collision scenarios and traffic violations to comprehensively assess safety.
-In the context of adversarial attacks, a higher infraction score is crucial for the safety of both passengers and pedestrians.
-As shown in Table IV, our model, P-CSG, demonstrates superior performance in two distinct adversarial scenarios: high-intensity, per-frame FGSM Attack, and uniformly applied, low-intensity Dot Attack.
+As shown in Table 4, P-CSG demonstrates superior performance in two distinct adversarial scenarios: high-intensity, per-frame FGSM Attack, and uniformly applied, low-intensity Dot Attack.
+Attack Model Driving score Route compl. Infrac. score
+[…]
+FGSM Attack $(\epsilon=0.01)$ P-CSG(Ours) $\mathbf{12.80}\pm 1.35$ $21.70\pm 2.87$ $\mathbf{0.59}\pm 0.02$
+Transfuser+ $2.69\pm 0.73$ $42.44\pm 3.65$ $0.22\pm 0.06$
+Interfuser* $8.37\pm 1.06$ $\mathbf{72.04}\pm 4.15$ $0.15\pm 0.04$
+Dot Attack P-CSG(Ours) $\mathbf{36.60}\pm 3.17$ $\mathbf{75.53}\pm 6.43$ $\mathbf{0.56}\pm 0.07$
+Transfuser+ $0.53\pm 0.21$ $54.06\pm 5.48$ $0.02\pm 0.01$
+Interfuser* $12.56\pm 2.15$ $73.62\pm 7.32$ $0.19\pm 0.02$
+Table 4: Attack Results.
 Specifically, P-CSG achieved an infraction score that was at least $2.5$ times higher than that of competing models in both attack scenarios tested.
-Under the low-intensity Dot Attack, P-CSG not only maintained high safety standards, indicated by a higher Infraction Score but also completed more routes, as reflected in a higher Route Completion metric.
+Under the low-intensity Dot Attack, P-CSG not only maintained high safety standards, indicated by a higher infraction score, but also completed more routes, as reflected in a higher route completion metric.
 Conversely, in scenarios involving high-intensity FGSM Attacks, P-CSG implemented a more conservative strategy aimed at enhancing security.
-This is shown by its higher Infraction Score, even though it led to fewer completed routes.
+Its higher infraction score shows this, even though it led to fewer completed routes.
 Most importantly, P-CSG consistently reaches the highest cumulative driving score in both types of attacks, confirming its robustness and adaptability in different challenging conditions.
-VI Conclusion
-In this paper, we improve the fusion technologies and policy learning approaches based on penalties for autonomous driving.
-Additionally, we compare our proposed model with other baselines under Dot Attacks and FGSM Attacks, demonstrating outstanding performance and robustness.
-We aspire for our proposed penalty-based imitation learning approach to introduce a fresh perspective into the domain of end-to-end autonomous driving, aimed at enhancing the compliance of autonomous agents with traffic rules.
-Furthermore, additional research is necessary to enhance the robustness of autonomous driving systems, particularly in countering attacks such as Dot Attacks.
+6 Conclusion
+In this paper, we improve the multi-modality fusion technologies and policy learning methods based on penalties for autonomous driving.
+We observe that our contrastive learning-based multi-modal fusion method is helpful in extracting the global content of various modalities.
+Furthermore, we found that incorporating penalties based on traffic rules into an imitation learning pipeline can enhance the agent’s adherence to those traffic rules.
+We also compare our proposed model with other baselines under Dot Attacks and FGSM Attacks.
+The results show that our approach is more robust against these attacks.
+We aspire for our proposed penalty-based imitation learning approach to introduce a fresh perspective into the domain of end-to-end autonomous driving to enhance autonomous agents’ compliance with traffic rules.
+One of our main objectives for future works is to include more types of penalties in the learning pipeline.
+These penalties could be designed based on multiple other traffic rules like speed limit, lane changing, and following distance.
+Scaling our method with a larger neural network and dataset is another direction of investigation, aimed at determining whether a larger model and data can effectively enhance overall performance.